### v2

# Through the Magnifying Glass: Adaptive Perception Magnification for Hallucination-Free VLM Decoding

Abstract
Existing vision-language models (VLMs) often suffer from visual hallucination, where the generated responses contain inaccuracies that are not grounded in the visual input. Efforts to address this issue without model finetuning primarily mitigate hallucination by reducing language biases reducing bias contrastively or amplifying the weights of visual embedding during decoding. However, these approaches improve visual perception at the cost of impairing the language reasoning capability. In this work, we propose the Perception Magnifier (PM), a novel visual decoding method that iteratively isolates relevant visual tokens based on attention and magnifies the corresponding regions, spurring the model to concentrate on fine-grained visual details during decoding. Specifically, by magnifying critical regions while preserving the structural and contextual information at each decoding step, PM allows the VLM to enhance its scrutiny of the visual input, hence producing more accurate and faithful responses. Extensive experimental results demonstrate that PM not only achieves superior hallucination mitigation but also enhances language generation while preserving strong reasoning capabilities. Code is available at https://github.com/ShunqiM/PM.
1 Introduction
The evolution of large vision-language models (VLMs) [40, 67, 61, 2, 70] has revolutionized the integration of computer vision and natural language processing, enabling seamless interaction between visual and textual modalities. These models effectively translate complex visual inputs into coherent and contextually relevant textual descriptions, driving advancements across a wide range of applications [65, 17, 60]. Fueled by innovations in model architectures [14, 18], training methodologies [50, 40, 12], and diverse multimodal datasets [53, 54], large VLMs continue to expand their capabilities and adaptability, solidifying their role as foundational tools in modern AI systems.
Despite their outstanding abilities, VLMs are prone to hallucination, where generated outputs fail to align with the visual input, with inaccuracies or unfaithful content. While some hallucinations stem from VLMs adopting the same architecture as pretrained large language models (LLMs) and inheriting their intrinsic limitations [42], the integration of visual and language features can also lead to over-reliance on language priors or limited visual representation capacity.
Figure 1: We propose Perception Magnifier (PM), a hallucination-mitigating decoding framework that locates and magnifies key visual regions based on inference attention patterns, enabling feature processing with finer details. For instance, PM locates the attention pattern on the top-right corner, and magnifies the glass bottle to it more discernible to the VLM.
Various approaches have been proposed to mitigate hallucination, ranging from training-end techniques, such as debiased dataset [39, 77] or enlarged visual resolution [75, 38], to inference end methods such as post-processing for hallucination removal [72, 82] and specialized decoding techniques [36, 24, 34]. Among these, hallucination-mitigation decoding methods have gained significant attention for their ability to improve visual faithfulness without requiring additional finetuning of the model. These methods primarily aim to enhance the contribution of visual information during model reasoning to mitigate hallucination. This enhancement can be achieved explicitly by increasing the weight of visual embeddings [43, 84], or implicitly by contrastively penalizing outputs that exhibit vague [34, 33], absent [79, 19], or confused visual interpretations [62, 5]. However, these methods face limitations in scenarios where insufficient visual resolution causes object misinterpretation. While a concurrent study [76] explores resizing a cropped key visual region as an additional input, it erases structural and contextual information from the image and confuses the VLM by giving multiple input images.
Inspired by the observation that VLMs allocate only marginal attention to many visual regions during inference, we propose the Perception Magnifier (PM), a visual decoding approach that adaptively modulates spatial resolution, processing key visual regions at higher detail while compressing less relevant areas, as illustrated in Fig. 1. By refining the visual input during decoding, PM enhances attention to fine-grained details without compromising structural visual context of the image or the intrinsic reasoning ability of VLM. Unlike methods that apply static modifications to input images [73, 68, 69, 76], PM dynamically adjusts the image representation throughout decoding, enabling more effective open-ended generation.
Specifically, we introduce a gradient-free approach to iteratively identify high-attention visual tokens relevant to generation across deep layers and aggregate their attention scores into a token-level heatmap. This heatmap is post-processed to enhance coverage and smoothness, then upsampled to construct a pixel-level perception map. Guided by this map, PM refines the visual input by magnifying relevant regions while compressing less critical areas, implemented as a resampling process [80, 27]. This dynamic adjustment allows the model to iteratively refine its focus throughout decoding, ensuring improved visual faithfulness without disrupting language reasoning.
PM is broadly applicable to interleaved VLM with unified visual sampling. Experiments on hallucination benchmarks along with GPT-4o-assisted assessments in both reference-based and vision-based settings, demonstrates that PM not only shows superior hallucination mitigation performance at question answering and open-end generation but also preserves the reasoning capabilities of VLMs. Additionally, we provide extensive ablations to analyze the contributions of each component in our method. Our contributions can be summarized as follows:
  • We propose the Perception Magnifier (PM), a decoding technique that mitigates hallucination by adaptively magnifying visual regions of interest during decoding.
  • We introduce a novel attention grounding method that iteratively aggregates attentions across layers to generate a perception map, which guides region magnification during decoding.
  • We conduct experiments with extensive evaluations, demonstrating the effectiveness of PM in mitigating hallucination while preserving reasoning capabilities.
2 Related Work
2.1 Vision-Language Models
The scaling up of computational resources and the availability of large-scale data has driven advances in pre-training [49, 15, 3] and instruction tuning techniques [48, 64, 45], allowing the emergence of numerous large language models (LLMs) [58, 81, 3, 46, 10] that can generate human-like text and perform complex reasoning across diverse tasks. Based on the advancements in LLMs, many vision-language models (VLMs) have been developed by integrating powerful language models with vision encoders using connection modules. While both cross-attention mechanisms [35, 12, 1, 83] and projector-based token concatenation [40, 2, 61, 59, 32, 60] were explored in earlier works, modern models [41, 22, 67, 6, 17] usually adopt the latter for its simplicity and efficiency in aligning visual and textual features. Although these VLMs leverage the power of existing LLMs by aligning image features with the text modality, their autoregressive nature and the limitations of visual model capabilities still lead to hallucination issues.
2.2 Hallucination in Vision-Language Models
In addition to hallucination inherited from the LLM backbones [42] due to insufficient context attention and capability misalignment [23, 78, 26], VLMs also face challenges unique to their multimodal nature, such as language bias [84, 79], where the language decoder may rely heavily on preceding textual tokens over visual tokens. To address this, various methods have been proposed to enhance the role of visual information by prompting input images [73, 68, 69] or amplifying the influence of visual tokens [79, 84, 43]. VLMs also hallucinate due to the limitations of vision encoders or connection modules [42]. Prior works have addressed this by increasing visual tokenization resolution [75, 38] or improving the capacity of these modules [41, 8]. Efforts to control hallucination have also focused on the data aspect, including mitigating bias [31, 39, 77, 56, 47, 74], creating biased samples for contrastive learning [28], and developing methods to evaluate hallucination [30, 21, 71, 63], supported by various metrics and algorithms for quantifying hallucination [52, 37, 44, 29, 7]. Post-processing techniques have also been proposed to eliminate hallucinations by identifying visually ungrounded outputs [72, 82]. While these approaches tackle hallucination from multiple angles, few methods focus on enhancing visual recognition capabilities beyond increasing resolution. Instead, we propose a novel approach that assists VLMs with a magnifier to capture fine-grained details in input images effectively. A concurrent but slightly earlier study, ViCrop [76], adopts a similar idea by cropping and resizing a key visual bounding box based on attention maps as additional inputs. However, it does destruct the integrity of image structure, and can fail when spatial awareness is required, whereas our method preserves the visual input structure by magnifying key regions while compressing less relevant areas.
2.3 Hallucination Mitigation Decoding Strategies
Decoding strategies determine how a model generates the next text token, with methods like beam search [57] enabling higher-quality text generation without additional fine-tuning. In the context of hallucination control, contrastive decoding (CD) [36] is a popular approach that optimizes output logits by amplifying patterns favored by larger expert LMs while suppressing a biased output more prone to hallucination, typically derived from smaller amateur LMs. DoLa [11] introduced an efficient CD formulation by treating early-exit logits of expert LMs as the biased output. Beyond CD, OPERA [24] mitigates hallucination in LMs by penalizing beam outputs that over-rely on specific aggregation tokens. For VLMs, VCD [34] extends CD by defining biased outputs as those generated with noisy image inputs. Subsequent works explored diverse biased output designs, such as logits derived without image inputs [79, 19], excluding key visual tokens [25], or producing hallucinative outputs via finetunning [5] or prompt-engineering [62]. Other methods aim to enhance expert outputs by increasing the weight of visual tokens in the model [43, 84] or searching for focal views with an external detector [9]. [33] further refines VCD by fusing outputs from multiple visually biased samples with an entropy-weighted strategy. Although these decoding methods effectively reduce hallucinations by penalizing less visually grounded outputs, treating them solely as ‘bias’ risks undermining the model’s ability to form logical inferences and maintain coherence in language-driven contexts. Instead, our method mitigates hallucination by adaptively magnifying relevant image regions, enhancing visual faithfulness while preserving the reasoning and language capabilities of VLMs.
3 Methods
Figure 2: Structural diagram of the proposed PM method. At each decoding step, we aggregate multi-layer attentions into a token-level heatmap $\mathcal{H}^{(i)}$. By iteratively masking the most attended tokens and re-forwarding, we collect heatmaps $\{\mathcal{H}^{(i)}\}$ and combine them into $\mathcal{H^{*}}$ once the process ends, and post-processing $\mathcal{H^{*}}$ to obtain $\mathcal{H}_{proc}$ (yellow block). The magnifier resamples the visual input according to $\mathcal{H}_{proc}$, producing $\hat{I}$ (purple block). The VLM can then generate the next token with enhanced focus on critical regions with finer details.
In traditional autoregressive VLMs, the model takes an image $I$ and a text sequence $T^{i}$ at each decoding step. Given this input, it predicts the next token $T^{i+1}$ with: $O^{i+1}=VLM(I,T^{i}),$ where $O^{i+1}$ represents the output logits over the vocabulary. The model then applies the softmax function to obtain a probability distribution: $P(T^{i+1}\mid I,T^{i})=\text{softmax}(O^{i+1}).$ A sampling strategy is then applied to determine the final token. For example, in greedy decoding, the token with the highest probability is selected: $T^{i+1}=\arg\max P(T^{i+1}\mid I,T^{i}).$ This predicted token is appended to the sequence, and the process repeats until the end-of-sequence (EOS) token is generated.
In this paper, we propose the Perception Magnifier (PM), a novel approach that iteratively identifies and enhances key visual regions to refine visual input for more effective token generation. As illustrated in Fig. 2, PM iteratively identifies highly attended visual tokens, blocks them, and re-forwards the sequence until no tokens receive significant attention. The identified tokens are then aggregated based on their attention scores to construct a perception map $\mathcal{P}$. This map is used to magnify the original image $I$ via a sampling process, $\hat{I}^{i}=\mathcal{T}(I,\mathcal{P})$. The refined image $\hat{I}^{i}$ then serves as the refined visual input for generating the next token, $T^{i+1}=\text{VLM}(\hat{I}^{i},T^{i})$, at the current decoding step. The following sections detail the perception map construction, iterative refinement for improving the coverage of perception map, and the attention-based magnification process in the proposed PM method.
3.1 Perception Map Construction
In this section, we first introduce a simple method to construct a naive perception map to guide the magnifier. This method is applicable to any mainstream VLMs that adopt an interleaved architecture, where patchified visual embeddings are integrated into the input sequence alongside textual tokens. These models process information by interleaving system prompts, visual representations, and past textual interactions, such as user queries and generated responses, into a unified sequence for a backbone LM.
Once the input sequence is formed, the transformer processes it via self-attention. For a single attention head, the attention weight from token $x_{i}$ to $x_{j}$ is computed as $\mathrm{Attn}_{i,:}=\text{Softmax}\left(Q_{i}K_{j}^{\top}/\sqrt{d_{k}}\right)$, where $Q_{i}$ and $K_{j}$ are the query and key vectors of tokens $x_{i}$ and $x_{j}$, respectively, and $d_{k}$ is the key dimension. The corresponding value $V_{j}$ is then weighted by $\mathrm{Attn}_{i,j}$ to obtain the hidden representation of $x_{i}$: $z_{i}=\sum_{j}\mathrm{Attn}_{i,j}\,V_{j}$. In practice, transformers stack multiple layers and heads, producing attention weights across different levels. These attention patterns can be represented as a multi-dimensional tensor: $\mathrm{Attn}\in\mathbb{R}^{N_{l},N_{h},H,W}$, where $N_{l}$ and $N_{h}$ denote the number of layers and attention heads, respectively, and $H\times W$ represents the spatial dimensions of visual tokens, which can be reshaped into a 2D feature map. PM utilizes the cached $\mathrm{Attn}$ to construct a token-level heatmap, highlighting the most relevant visual regions based on the model attention.
As indicated by prior studies [13], deep networks primarily aggregate and interpret visual information in their earlier layers before leveraging these representations in later layers. Moreover, it has been observed that intermediate features can already yield competitive performance [11]. In line with these findings, we empirically note that middle-layer attention often more accurately localizes relevant objects than that of the final layers. Hence, we construct the perception map by aggregating self-attention matrices from layers $l\geq\mathcal{L}$ and applying a max pooling across all heads $h$ to highlight the impact of more visually important regions. Formally, the token-level heatmap $\mathcal{H}$ can be obtained by:
$\mathcal{H}=\sum_{l=\mathcal{L}}^{N_{l}}\max_{h\in{1,\ldots,N_{h}}}\mathrm{Attn}_{l,h}.$ (1)
Such $\mathcal{H}\in\mathbb{R}^{H,W}$ reflect the intermediate reasoning of VLM and highlight the salient regions identified.
Post-Processing to Pixel Space.
The naive perception map constructed without iterative refinement could already be upsampled to guide the magnifier. Before up-sampling the token-level heatmap $\mathcal{H}$ to pixel-level perception map, we first post-process the heatmap with the method similar to [73] to obtain a more smooth and well-distributed attention distribution for better magnification quality. Firstly, we normalize $\mathcal{H}$ by scaling them to unit interval, obtaining $\mathcal{H}^{\prime}$. Then, we amplify the variance of the normalized map by a scaling coefficient of $\alpha$, and then compress its values back into the range $[0,1]$ using a sigmoid function, resulting in $\mathcal{H}_{\text{proc}}$:
$\mathcal{H}_{\text{proc}}=\text{sigmoid}\!\left(\alpha\cdot\,\frac{\mathcal{H}^{\prime}-\mu(\mathcal{H}^{\prime})}{\sigma(\mathcal{H}^{\prime})}\right).$ (2)
Finally, we apply a uniform smoothing filter with a kernel size of $k$ to the enhanced heatmap, followed by bilinear upsampling to obtain the pixel-level perception map $\mathcal{P}$. As shown in Fig. 2, this post-processing step refines the heatmap by amplifying low-magnitude attention values, ensuring that small but potentially important regions are also considered for magnification.
3.2 Iterative Refinement
While the naive perception map serves as a simple and effective guide for magnification, a single forward pass may fail to capture all instances of objects when multiple occurrences exist. This limitation arises because deep vision models exchange fine-grained features at lower layers and aggregate them into specific tokens that act as information registers [13, 51]. This aggregation concentrates visual information into a few key tokens, reducing the contribution of other visual tokens associated with semantically similar regions. As a result, when similar visual patterns appear in different parts of the image, VLMs may primarily attend to a subset of key tokens, leading to an incomplete perception map that overlooks certain regions of interest.
To enhance the coverage of perception map, we propose an iterative refinement approach on attention that compels VLMs to attend to additional visual regions by iteratively blocking key registers from their reasoning process. However, directly masking key visual regions at the image level would require fine-grained segmentation or ROI localization, making it computationally challenging. Moreover, such an approach is infeasible for decoding, as it would necessitate re-tokenization of visual tokens. Instead, we apply information blocking at the token level, effectively mitigating these challenges while preserving inference efficiency.
Specifically, at the first iteration, we forward the tokenized input through the VLM using the default attention mask $M^{(0)}$, initialized as a vector of ones, and obtain the initial token-level attention heatmap $\mathcal{H}^{(0)}$. To identify the most visually attended tokens $T^{(0)}$, we employ a clustering-based method that partitions the attention values into two clusters, selecting tokens in the one with the higher centroid as $T^{(0)}$, as visually important tokens typically exhibit significantly higher attention than others. To suppress their influence in subsequent iterations, we update the attention mask as: $M^{(1)}=M^{(0)}\cap\neg T^{(0)}$, where $\neg T^{(0)}$ represents the complement of the mask over $T^{(0)}$, ensuring that these highly attended tokens are excluded in the next pass. The updated input is then reprocessed by the VLM, producing a new token-level heatmap $\mathcal{H}^{(1)}$. This process is repeated iteratively until the total attention score in the next iteration falls below $\beta$ or an upper limits on iterations is reached, i.e., $\sum\mathcal{H}^{(i+1)}<\beta,\quad\text{for }0\leq i\leq 8,$ indicating that no additional visually relevant tokens remain to be allocated substantial amount of attention.
Finally, we construct the refined token-level heatmap $\mathcal{H}^{*}$ by aggregating attention across all $T$ iterations while normalizing for repeated contributions at each token position:
$\mathcal{H}^{*}=\frac{\sum_{i=0}^{T}\mathcal{H}^{(i)}}{\sum_{i=0}^{T}\mathbbm{1}_{M^{(i)}}}.$ (3)
The iteratively refined heatmap $\mathcal{H}^{*}$ is then post-processed to $\mathcal{H}_{proc}$ using the aforementioned method to produce a pixel-level perception map $\mathcal{P}$. This refined perception map captures all major visual elements relevant to next-token generation, providing an enhanced guide for the perception magnifier and ensuring focus on the most critical regions for the VLM.
3.3 Attention-Based Magnification
Inspired by [80], we develop the perception magnifier as an attention-based sampling process. The perception map $\mathcal{P}$ is treated as a mass function, where regions with higher perception values are more likely to be sampled. To achieve this, we first decompose $\mathcal{P}$ into marginal distributions along the horizontal and vertical dimensions and compute their cumulative distributions $\mathcal{F}_{x}(n)$ and $\mathcal{F}_{y}(n)$ as follows:
$\begin{split}\mathcal{F}_{x}(n):=\sum_{j=1}^{n}\max_{1\leq i\leq w}\mathcal{P}_{i,j},\\ \mathcal{F}_{y}(n):=\sum_{i=1}^{n}\max_{1\leq j\leq h}\mathcal{P}_{i,j},\end{split}$ (4)
where $w$ and $h$ denote the width and height of the perception map, respectively. Next, we apply inverse transform sampling to remap the original image coordinates to resampled coordinates, ensuring structure-preserving magnification along both dimensions. The magnified image is then constructed as:
$\hat{I}=\mathcal{T}(I,\mathcal{P})_{i,j}=\text{Interp}\big{(}I,\mathcal{F}_{x}^{-1}(i),\mathcal{F}_{y}^{-1}(j)\big{)},$ (5)
where $\mathcal{F}^{-1}(\cdot)$ denotes the inverse function of $\mathcal{F}(\cdot)$, which remaps uniformly spaced output coordinates to a perception-aware warped space. To ensure smooth spatial mapping, we use bilinear interpolation as the Interp method in the equation, balancing efficiency and effectiveness, though other interpolation methods could also be applied. The magnified image $\hat{I}$ replaces the original visual input $I$ and is forwarded through the VLM with the same textual input to obtain updated model output logits $\hat{O}$. While perception-magnified logits $\hat{O}$ can be proportionally combined or contrastively compared with the original model output $O$ for next-token generation, we directly sample from $\hat{O}$ after the softmax function, as it already provides an enhanced representation of the visual content. The entire perception magnification process is applied at each auto-regressive decoding step, enabling the model to adaptively focus on different regions of the visual input while preserving structural integrity, hence enhancing perception ability and mitigating hallucination.
4 Experiments and Results
Object-level Attribute-level MME Score POPE Accuracy
Existence Count Position Color Total* AOKVQA GQA COCO AVG*
Greedy 195.00 143.33 128.33 163.33 630.00 84.10 84.39 85.29 84.59
OPERA [24] 195.00 143.33 128.33 163.33 630.00 84.09 84.43 85.29 84.60
VCD [34] 183.33 143.33 110.00 150.00 586.67 84.18 82.32 85.49 83.99
ICD [57] 190.00 148.33 123.33 153.33 615.00 84.23 84.89 84.92 84.68
SID [25] 195.00 155.00 126.67 160.00 636.67 84.39 84.17 85.81 84.79
VDD [79] 190.00 140.00 133.33 165.00 628.33 86.73 85.52 86.71 86.32
PAI [43] 190.00 155.00 133.33 170.00 648.33 85.32 83.43 86.66 85.14
IBD [84] 190.00 160.00 133.33 170.00 653.33 86.37 84.80 86.65 85.94
API-C [73] 190.00 150.00 128.33 165.00 633.33 86.87 85.04 87.31 86.41
API-L [73] 190.00 150.00 133.33 168.33 641.67 86.52 84.30 86.60 85.81
ViCrop-R [76] 190.00 163.33 105.00 175.00 633.33 85.53 85.19 87.43 86.05
ViCrop-G [76] 195.00 153.33 96.67 175.00 620.00 86.09 85.18 87.31 86.19
PM (Ours) 195.00 175.00 138.33 175.00 683.33 86.94 85.46 87.68 86.70
Table 1: Evaluation results for MME Perception Scores (left) and POPE Accuracy (right). Columns marked with * indicate the overall scores for the benchmark. The result for BEAM is omitted for its equivalence with OPERA on single-token QA. API and ViCrop each propose multiple variants, denoted by postfixes (e.g., -C or -L), with details provided in the Supplementary Material. The proposed PM method consistently demonstrates superior performance on both benchmarks.
4.1 Experimental Settings
Datasets and Baselines.
We evaluate our method on three benchmarks: MME [20], POPE [37], and LLaVA-Bench [40], all commonly used for assessing multimodal hallucinations. For comparison, we include various decoding strategies, including greedy decoding, contrastive decoding-based methods (VCD [34], VDD [79], IBD [84], PAI [43], ICD [62], SID [25]), beam search [57], OPERA [24], API [73], and ViCrop [76]. Further details are provided in the Supplementary Material.
Implementation Details.
We conduct our experiments using LLaVA-1.5 7B [41], a widely adopted VLM known for its strong visual interpretation and language understanding capabilities, as well as its computational efficiency. For hyperparameters, we start aggregating attentions from layer $\mathcal{L}=12$, set the scaling coefficient and kernel size to $\alpha=10$, $k=3$, and set the attention threshold for iterative refinement to $\beta=0.3$.
4.2 Quantitative Analysis
Tab. 1 presents the evaluation results of various methods on both the MME and POPE benchmarks. Our proposed method consistently outperforms all existing approaches by a substantial margin. Compared to MME, POPE primarily focuses on existence-based questions, allowing most methods to improve performance by reinforcing the contribution of visual tokens in model inference, thereby enhancing perception ability. In contrast, even within the perception subset, MME poses greater challenges by incorporating more complex questions that require both fine-grained visual recognition and reasoning.
Empirically, CD-based approaches operate at the logit level by contrasting visually intact or compromised outputs. While this strategy can enhance perception by suppressing hallucinated content, it fails to resolve cases where fine-grained details remain unrecognized across both logits. Methods that enhance visual representation at the embedding level, such as PAI and IBD, are less affected by this limitation, as they reinforce visual contributions at multiple layers. This enhancement may implicitly emphasize fine-grained details during intermediate processing, but it can also disrupt the balance between visual and linguistic reasoning, leading to potential trade-offs. By modifying input images, API strengthens visual contributions without altering the embedding distribution of VLM, but it primarily mitigates the impact of less relevant regions by applying a grayscale mask. While this reduces interference from irrelevant areas, it does not explicitly refine fine-grained details, limiting its enhancement on the perception of VLMs on subtle visual cues. ViCrop also exhibits limitations on MME, as isolating a region through cropping removes its surrounding visual context, making position-related questions more challenging. Additionally, providing the cropped image as an additional input can introduce ambiguity, hindering the VLM’s ability to interpret spatial relationships and maintain a coherent understanding of the scene.
In contrast, our method enhances critical visual information at the input level while preserving the VLM’s internal state, maintaining a stable embedding distribution across layers. It further improves fine-grained recognition by adaptively magnifying relevant regions while preserving contextual information. As a result, our approach effectively reduces hallucinations and consistently outperforms existing methods on benchmarks.
4.3 GPT-4o Assisted Evaluation
LLaVA-Bench COCO In-the-Wild
Ref $\uparrow$ Acc $\uparrow$ Det $\uparrow$ Ref $\uparrow$ Acc $\uparrow$ Det $\uparrow$
Greedy 81.15 92.69 95.76 56.94 71.83 79.95
BEAM 83.17 94.75 91.60 58.94 74.71 78.25
OPERA [24] 82.99 96.12 93.61 58.44 71.81 77.92
VCD [34] 82.00 92.21 94.39 57.17 69.4 81.22
ICD [57] 83.29 93.91 95.46 54.43 66.89 74.81
SID [25] 81.10 97.35 95.80 59.49 70.93 81.63
VDD [79] 81.38 95.62 102.75 61.60 78.22 83.01
PAI [43] 79.04 92.06 90.59 52.29 69.74 72.49
IBD [84] 85.29 95.73 95.45 61.59 77.91 82.86
PM (Ours) 85.76 97.74 95.80 62.46 77.21 84.50
Table 2: Results on LLaVA-Bench (%). Ref denotes reference-based evaluation, while Acc and Det represent accuracy and detailedness in vision-based evaluation, respectively.
Previous research [24, 84] has shown that GPT-4-assisted evaluation is a reliable metric for assessing hallucination. Accordingly, we adopt both reference-based and vision-based evaluations on LLaVA-Bench, using GPT-4o as the evaluator. In reference-based evaluation, GPT-4o receives the question with a detailed textual description of the image, while in vision-based evaluation, it is provided with both the question and the image.
Under both settings, GPT-4o scores the answer generated by the tested method and a reference answer from GPT-4 on a scale of 0 to 10. The final score is computed as the ratio of the tested method’s score to GPT-4’s score. For reference-based evaluation, GPT-4o focuses on overall content quality, whereas for vision-based evaluation, it examines the image and assesses accuracy and detailedness. We exclude API [73] and ViCrop [76] from these benchmarks, as they are not designed for open-end generation. Complete evaluation prompts are provided in the Supplementary.
As shown in Tab. 2, the proposed PM achieves the most competitive overall performance on open-end generation tasks. While its detailedness on COCO and accuracy on the In-the-Wild split are slightly lower than a few existing method, PM still demonstrates advantage in hallucination mitigation, particularly in reference-based evaluation. These results indicate that even in open-end decoding where the model does not always need to focus on specific regions at each step, the proposed adaptive perception magnifier substantially enhances output quality, demonstrating its effectiveness in reducing hallucinations.
4.4 Maintaining Cognition Ability
MME-Cognition CS NR Trans CR Total*
Greedy 112.86 70.00 107.50 67.50 357.86
BEAM 112.86 47.50 90.00 60.00 310.36
OPERA [24] 112.86 47.50 90.00 60.00 310.36
VCD [34] 120.00 37.50 97.50 72.50 327.50
ICD [62] 120.71 70.00 92.50 65.00 348.21
SID [25] 122.86 55.00 85.00 65.00 327.86
VDD [79] 110.14 58.00 68.00 86.00 322.64
PAI [43] 116.43 70.00 100.00 67.50 353.93
IBD [84] 117.14 70.00 107.50 57.50 352.14
API-C [73] 115.00 47.50 107.50 47.50 317.50
API-L [73] 120.00 45.00 115.00 45.00 325.00
ViCrop-R [76] 115.00 50.00 105.00 65.00 335.00
ViCrop-G [76] 109.29 50.00 105.00 86.00 350.29
PM (Ours) 117.86 42.50 107.50 87.50 355.36
Table 3: Comparison of different hallucination control decoding methods on the MME-Cognition subset, including Commonsense Reasoning (CS), Numerical Reasoning (NR), Translation (TR), and Code Reasoning (CR).
Many hallucination control decoding methods enhance perception by increasing the contribution of visual information during token generation. While these approaches generally improve perceptual ability, they often disrupt the calibrated balance between the vision and language modalities, potentially compromising the inherent cognitive capabilities learned by the VLM during large-scale pretraining.
To investigate this, we evaluate our method and baselines on the Cognition split of MME, where answering requires reasoning beyond direct visual cues. As shown in Tab. 3, most CD-based methods that explicitly modify output logits exhibit a notable decline in cognition ability, aligning with expectations. Surprisingly, PAI and IBD, which enhance the attention weight of visual tokens through intermediate embedding modifications, do not incur a substantial loss in cognition performance. This suggests that the bias introduced by their layer-wise manipulation is relatively mild and can be compensated for by subsequent model layers. Furthermore, despite API-based methods modifies the image without altering the VLM’s internal reasoning, they still lead to degraded performance. We hypothesize that the gray masking applied to images in the cognition set where backgrounds are typically uniform may introduce additional perceptual burden, ultimately impacting overall performance.
In contrast, our proposed PM method does not negatively affect cognition. Although our approach is designed solely for hallucination control rather than cognition enhancement, the total cognition score remains comparable to greedy decoding and outperforms other methods, demonstrating its robustness in preserving reasoning capabilities.
4.5 Qualitative Analysis
Question:
Is there a chair
in the image?
Greedy: No
PM: Yes
Question:
Is there a tree
in the image?
Greedy: Yes
PM: No
Figure 3: Qualitative examples on the POPE benchmark showing the input question, the original image, the processed perception map $\mathcal{H}_{\text{proc}}$, and the perception-magnified image, arranged from left to right. The answers generated by greedy decoding and PM are compared, with correct answers highlighted in green and incorrect ones in red.
Figure 3 presents qualitative examples illustrating the input question, the original image, the processed perception map $\mathcal{H}_{\text{proc}}$, and the perception-magnified image. The results demonstrate that the proposed method effectively enlarges small objects relevant to the question while preserving fine details. In the first example, the VLM initially attends to the top-left region, recognizing the potential presence of a chair but failing to identify it. With the magnified input, the chair is presented at a higher spatial resolution, enabling the VLM to correctly recognize it. Similarly, in another case, the VLM misidentifies an object behind a door window as a tree. However, the perception-magnified input allows the model to re-examine the object with greater detail, preventing hallucinated predictions.
Overall, the qualitative results suggest that the proposed method enhances the perceptibility of fine-grained visual details in regions critical for next-token generation. By enabling the model to examine these areas at a higher resolution, our approach effectively reduces hallucinations and improves overall model performance.
4.6 Ablations
MME-Perception Object-level Attribute-level
Existence count position color Total
Greedy 195.00 143.33 128.33 163.33 630.00
PM w/o IR & MLA 195.00 155.00 120.00 170.00 640.00
PM w/o MLA 195.00 155.00 125.00 170.00 645.00
PM w/o IR 190.00 175.00 135.00 165.00 665.00
PM 195.00 175.00 138.33 175.00 683.33
Table 4: Ablations on iterative refinement (IR) and multi-layer aggregation (MLA). Both components contribute to performance improvements compared to their respective baselines.
Iterative Refinement.
We ablate the impact of iterative refinement in constructing the perception map for the magnifier. As shown in Tab. 4, the refined perception maps consistently outperforms their counterpart without iterative refinement, demonstrating the effectiveness of our approach. These results indicate that iterative refinement alleviates the effects of information aggregation during perception map construction, leading to more precise coverage of visually relevant regions for next-token generation. While iterative refinement substantially enhances the accuracy of the perception map, it does not entirely prevent information leakage from masked tokens. This limitation arises from the inherent feature mixing among visual tokens during early-stage processing, as most VLMs employ a Vision Transformer [16] for feature extraction. Despite this, our method achieves comprehensive coverage of key visual regions while maintaining a strong balance between attention precision and computational efficiency, making it a robust and effective solution for guiding the magnifier.
MME-Perception Object-level Attribute-level
Existence Count Position Color Total
Greedy 195.00 143.33 128.33 163.33 630.00
API CLIP [73] 190.00 155.00 131.67 170.00 646.67
API LLaVA [73] 190.00 165.00 126.67 170.00 651.67
ViCrop-R [76] 195.00 163.33 116.67 170.00 645.00
ViCrop-G [76] 195.00 158.33 116.67 170.00 640.00
Rel Map [4] 195.00 175.00 130.00 155.00 655.00
PM w/o IR 190.00 175.00 135.00 165.00 665.00
Table 5: Comparison of different perception maps for guiding the magnifier. IR denotes iterative refinement. Our perception map construction method achieves superior results even without IR.
MME-Perception Object-level Attribute-level
Existence count position color Total
Greedy 195.00 143.33 128.33 163.33 630.00
Blurring 185.00 158.33 111.67 175.00 630.00
Bounding Box 180.00 160.00 130.00 170.00 640.00
Crop 195.00 155.00 128.33 170.00 648.33
Crop & Resize 195.00 153.33 123.33 175.00 646.67
Magnification 195.00 175.00 138.33 175.00 683.33
Table 6: Comparison of methods for highlighting key visual regions based on the perception map.
Multi-Layer Aggregation.
We also ablate the impact of multi-layer aggregation in constructing the perception map for the magnifier. As shown in Tab. 4, perception maps generated by aggregating attention patterns across multiple layers exhibit substantial differences from those constructed using attention from a single layer. Empirically, while VLMs tend to focus more on visual features in middle-to-deep layers, the specific layer receiving the strongest visual attention varies unpredictably. This observation suggests that relying on a single layer may lead to inconsistencies in capturing visual information. Our analysis shows that in most cases, a random subset of intermediate layers exhibits a clear and structured visual attention distribution, whereas in other layers, the model allocates little to no attention to visual tokens, apart from a few sink tokens [66]. This variability highlights the necessity of aggregating attention patterns across multiple layers to ensure that the most attended visual features across layers are effectively incorporated into the perception map.
4.7 Quality of Perception Map
Existing studies [4, 55] use attention-based relevancy maps to improve VLM interpretability by aggregating attention weights with backpropagated gradients. API [73] constructs attribution maps using either CLIP’s CLS token decomposition or averaged attention weights in LLaVA, while ViCrop [76] introduces attention aggregation techniques that vary in their reliance on gradient-based information. Although developed for different purposes, these methods can also be adapted for perception map construction.
Building on these methods, we compare their heatmap quality against our perception map, using the base version from Sec. 3.1 without iterative refinement and applying identical post-processing. As shown in Tab. 5, our approach achieves superior perception quality. CLIP-based maps prioritize global image-text similarity rather than token-level relevance, while API-LLaVA and ViCrop-R aggregate attention from a single intermediate layer and average across heads, leading to reduced stability and diluted key features. In contrast, our max-pooling approach better retains critical visual information. The Rel map method and ViCrop-G employ backpropagation-based fine-grained attributions but provide limited benefits for magnification, as they focus on token-level attribution with more precise attention values but without necessarily improving coverage. Furthermore, their reliance on computationally expensive gradient updates makes them less practical for efficient perception-guided magnification.
4.8 Image Prompting Method
Various methods have been proposed to “prompt” images to emphasize specific regions [69, 76], including add a bounding box on the highlighted region, or apply blurring effect, crop the remaining regions, or crop and resize the key area. In this section, we benchmark these methods as alternatives to our magnifier, evaluating their effectiveness in highlighting key visual regions for focus-adaptive decoding when combined with the constructed perception map. As shown in Tab. 6, our magnification approach effectively leverages the perception map to enhance the perception ability of the VLM, whereas other prompting techniques provide only marginal improvements over greedy decoding.
This difference arises because magnification not only highlights important regions but also reveals additional fine-grained details within limited resolution, allowing the model to capture subtle visual patterns. In contrast, many prompting methods merely emphasize region importance without increasing effective resolution, limiting their impact on fine-grained recognition. Meanwhile, when combined with ViCrop, the proposed method still improves performance, as ViCrop also enhances local details through cropping and resizing. However, ViCrop discards contextual information from the image, leading to reduced effectiveness in tasks that require spatial awareness, such as position-related questions.
5 Conclusion
We introduce the Perception Magnifier (PM), a visual decoding method that refines image inputs to mitigate hallucinations in vision-language models. PM adaptively enlarges visually critical regions identified by multi-layer attention iteratively, providing enhanced effective resolution on relevant finer-grained details at each decoding step. Experiments on MME, POPE, and LLaVA-Bench demonstrate that PM significantly reduces hallucinations at QA and open-end generation, while preserving reasoning capabilities. Future work may focus on optimizing the decision process for input magnification at each step, refining the identification of salient tokens, and extending the magnification-based approach to a broader range of VLM architectures.
Supplementary Material
Object-level Attribute-level MME Score POPE Accuracy
Existence Count Position Color Total* AOKVQA GQA COCO AVG*
Greedy 190.00 160.00 71.67 155.00 576.67 81.87 81.27 83.46 82.19
VCD [34] 185.00 163.33 65.00 148.33 561.67 81.86 81.14 83.31 82.10
ICD [57] 195.00 165.00 65.00 145.00 575.00 81.93 81.05 83.78 82.25
SID [25] 195.00 155.00 73.33 145.00 568.33 81.57 81.40 83.43 82.13
VDD [79] 195.00 155.00 73.33 155.00 578.33 81.99 81.74 83.03 82.26
API-C [73] 185.00 165.00 76.67 165.00 591.67 82.21 81.28 83.47 82.32
API-L [73] 190.00 165.00 73.33 160.00 588.33 81.84 81.24 83.33 82.14
PM (Ours) 190.00 165.00 68.33 190.00 613.33 82.15 81.27 83.66 82.36
Table 7: Evaluation results for MME Perception Scores (left) and POPE Accuracy (right) on the Shikra VLM. Columns marked with * indicate the overall scores for the benchmark. Some methods are omitted due to limitations on computational resources and implementations. The proposed PM method consistently demonstrates superior performance on both benchmarks.
6 Datasets
We carry out our experiments on three benchmarks popular for examining the perception ability and hallucination of VLMs.
MME [20] benchmark is designed to comprehensively assess both perception and cognition abilities of VLMs across 14 distinct subtasks on 2,370 questions. In our study, we focus primarily on the Perception subset of MME following [34] to evaluate the effectiveness of our method in controlling hallucinations.
POPE (Polling-based Object Probing Evaluation) [37] is a 9 $\times$ 3,000 question benchmark for evaluating object hallucination in VLMs. It formulates hallucination evaluation as a binary classification task by posing yes/no questions about object presence, addressing limitations of CHAIR [52], which relies on exact text matching, and providing a more robust evaluation less sensitive to prompt variations.
LlaVA-Bench [40] is a benchmark designed to evaluate VLMs in real-world visual chat scenarios. It comprises two components: LLaVA-Bench (COCO), utilizing images from the COCO dataset, and LLaVA-Bench (In-the-Wild), a more challenging subset featuring 24 diverse images—such as indoor and outdoor scenes, memes, paintings, and sketches—paired with 60 detailed, manually curated questions. The benchmark covers various vision-language tasks, making it suitable for assessing model generalization. In our study, we employ both COCO and In-the-Wild splits to comprehensively evaluate different methods on hallucination control, cognitive abilities, and performance in open-ended question answering.
7 Baselines
This section summarizes the high-level idea of methods used as comparisons in the experiments. We apply these methods on our benchmarks with the default hyperparameters provided in the original paper.
Beam Search explores multiple output candidates by iteratively expanding the most probable sequences and selecting the top-$k$ hypotheses at each step, resulting in improved answer diversity but at the cost of increased computational complexity proportional to the beam size. We evaluated all beam-based methods with a beam size of 4.
OPERA [24] is a beam-search-based method, examining the presence of a hallucinate beam and pruning the beam based on its attention on information aggregation patterns.
VCD [34] introduces noise into the input image and processes it through the VLM to generate a hallucinative output. By leveraging contrastive decoding (CD) to penalize high-probability predictions that align with the hallucinative output, VCD reduces the likelihood of hallucinations in the final prediction, resulting in more reliable responses.
ICD [62] generates a hallucinative output for CD-based decoding by appending ”You are a confused object detector” to the input prompt.
SID [25] induces a hallucinative output for CD-based decoding by identifying key visual tokens in the image and pruning them through forward.
VDD [79] mitigates language-induced hallucinations in VLMs by generating a hallucinative output through the removal or replacement of visual tokens with <unk> tokens.
PAI [43] enhances visual faithfulness by increasing the weights of visual tokens through attention mechanisms while penalizing outputs generated without visual inputs using CD.
IBD [84] also biases the output toward visual inputs by adding visual attentions and refines it against the original output using CD. Additionally, it adaptively adjusts the contrastive update strength at each decoding step based on contextual needs.
API [73] highlights key visual regions by overlaying a gray-level mask on the input image. The mask is guided by a heatmap derived from image-text associations using a CLIP model (-C) or a LLaVA model (-L).
ViCrop [76] estimates a bounding box around the key region in the image and generates a new input by cropping and resizing this region. Three approaches are proposed for bounding box estimation with different level of reliance on gradient information: relative attention (-R), gradient attention (-G), and pure gradient (-P). The pure-gradient approach is not benchmarked due to its higher GPU memory consumption, which necessitates lower model precision and results in degraded performance, making it unsuitable for a fair comparison.
8 Performance on Shikra
To assess the generalizability of PM beyond LLaVA, we apply it, along with several comparative methods, to Shikra-7B [6], a VLM designed for referential dialogue. The performance is evaluated on MME and POPE using Shikra with iterative refinement threshold $\beta=1.5$. The results are presented in Tab. 7. Note that certain comparative methods are excluded either due to the lack of available implementations for Shikra or constraints on computational resources.
As shown in Tab. 7, the proposed method continues to improve performance on Shikra compared to the baselines, though the improvement is less pronounced than on LLaVA. This discrepancy may stem from Shikra’s lower visual resolution and fewer visual tokens, which impact the effectiveness of perception magnification. With fewer visual tokens, cases where an object is partially recognized at intermediate layers but not reflected in the final prediction become less frequent, as each token carries greater attention weight in the inference process. Therefore, the proposed method has fewer opportunities to enhance recognition. Additionally, the reduced number of visual tokens lowers the granularity of the token-level perception map, leading to less precise perception magnification, which may further limit the performance gains of PM.
This observation suggests that the proposed method could be particularly beneficial for high-resolution VLMs, where small local features are more likely to be detected at intermediate layers but fail to significantly influence the final prediction. However, verifying this hypothesis is constrained by computational resource limitations.
Overall, the results on Shikra provide strong evidence that the proposed PM method generalizes well beyond LLaVA and effectively enhances hallucination mitigation in applicable VLMs.
9 Qualitative Samples on PM
Question:
Is there a building
in the image?
Greedy: No
PM: Yes
Question:
Is there a bottle
in the image?
Greedy: No
PM: Yes
Question:
Is there a paper
in the image?
Greedy: Yes
PM: No
Question:
Is there a cabinet
in the image?
Greedy: Yes
PM: No
Figure 4: Additional qualitative examples on the POPE benchmark showing the input question, the original image, the processed perception map $\mathcal{H}_{\text{proc}}$, and the perception-magnified image, arranged from left to right. The answers generated by greedy decoding and PM are compared, with correct answers highlighted in green and incorrect ones in red.
Fig. 4 presents additional qualitative examples illustrating the input question, the original image, the processed perception map $\mathcal{H}_{\text{proc}}$, and the perception-magnified image. The results demonstrate that the proposed method effectively enlarges small objects relevant to the question while preserving fine details. In many cases, the VLM benefits from the perception-magnified image, as it can now recognize previously overlooked objects that were too small to be detected in the original image. This leads to more accurate predictions, as observed in the first and second rows.
Furthermore, when the VLM erroneously identifies a non-existent object as present, as shown in the third example, PM magnifies the corresponding region, enabling the model to reassess the scene with finer detail. This process helps reduce false positive hallucinations by magnifying the visual details hence improve perceptibility. However, localized magnification does not always guarantee improved visual faithfulness. As demonstrated in the fourth row, despite providing a closer view, the model now fails to recognize the cabinet in the background. This may be due to the cabinet being identified based on its structural context rather than fine-grained textures, highlighting the limitations of region-focused magnification.
Overall, these results suggest that the proposed method effectively enhances the perceptibility of visual detail on regions crucial for next-token generation while generally preserving the overall structural integrity of objects, ultimately reducing hallucinations and improving model performance.
10 Limitations
Figure 5: A sample CLEVR-style image (left) and its corresponding magnified visual input (right). For tasks involving shape or size-related attributes, the proposed PM may introduce unintended distortions due to unregulated magnification. In this example, magnifying both the red cube and the blue sphere results in an enlarged sphere and a deformed cube, which may negatively affecting model predictions.
While the proposed PM method improves VLM inference by mitigating hallucinations, magnifying image regions may also introduce certain drawbacks. This section discusses key limitations of our approach.
Shape Distortion in Region Magnification.
Although region magnification allows the model to observe fine-grained visual details without altering the overall image structure, it may distort local shape properties. Specifically, in tasks requiring precise shape or length recognition, such as distinguishing letters or assessing object sizes, PM does not guarantee proportional magnification. This can introduce challenges for accurate recognition and induces hallucinations. For instance, when applied to CLEVR-style questions, as shown in Fig. 5, with the query “Please estimate how many times large is the blue sphere compared to the red cube?”, PM magnifies both objects, causing the blue sphere to appear visually larger while distorting the shape of the other two cubes. Such unintended modifications may negatively impact final predictions. Future work could explore refinements in perception map construction or post-processing techniques to achieve more structure-preserving magnification.
Improving Compatibility.
While the proposed method can be easily applied to interleaved VLMs with visual tokens corresponding to all patch features, applying it to VLMs where the correlation between visual tokens and image regions are non-trivial, can be slightly more complicated. The method depends on certain mapping from the visual token attention weights to the attention span on the original image. Therefore, when apply PM to models where the visual tokens are constructed with not only a feature extractor, but also a token sampler or cross-attention feature extractors, PM will need some attention mapping mechanism between visual tokens and patches. However, we can resolve this problem by introducing gradient backpropagation to address the implicit relationships between the image and tokens, refining the visual attentions and construct the perception map, similar to method proposed in [4]. Overall, such limitation does not reduce the contribution of this work, as the idea of adaptive-magnification can also always be used in aligned with interpretability tools designed for different VLMs. Alternatively, one could use PM in ways similar to [73], where any compatible model could be used as external guidance to the decoding-adaptive magnification.
11 Prompts for GPT-Assisted Evaluation
We adopt the GPT-assisted evaluation prompts directly from prior studies. For reference-based evaluation, we follow the prompt used in [40], with an example of the prompt and corresponding evaluation result provided in Tab. 8. For vision-based evaluation, we use the prompt from [24], with an example shown in Tab. 9.
Prompt for Reference-based Evaluation [Context] An aerial view of Diamond Head in the Hawaiian Islands. [Question] What is the name of this famous sight in the photo? [Assistant 1] The famous sight in the photo is Diamond Head. [End of Assistant 1] [Assistant 2] The famous sight in the photo is the Na Pali Coast, which is a scenic coastal area on the island of Kaua’i in Hawaii. [End of Assistant 2] [System] We would like to request your feedback on the performance of two AI assistants in response to the user question displayed above. The user asks the question on observing an image. For your reference, the visual content in the image is represented with a few sentences describing the image. Please rate the helpfulness, relevance, accuracy, level of details of their responses. Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance. Please first output a single line containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgement. Response: 10 2 Assistant 1 correctly identified the famous sight in the photo as Diamond Head, which is consistent with the context provided. The response is accurate, relevant, and directly answers the question with the appropriate level of detail given the context. Assistant 2, on the other hand, incorrectly identified the sight as the Na Pali Coast, which is not mentioned in the context. This response is inaccurate and irrelevant to the question, as it does not align with the description of the image provided. Additionally, the details about the Na Pali Coast, while informative, are not applicable to the context of Diamond Head. Therefore, Assistant 2’s response is not helpful in this scenario.
Table 8: Prompt for reference-based evaluation.
Prompt for Vision-based Evaluation [Uncaptioned image] You are required to score the performance of two AI assistants in describing a given image. You should pay extra attention to the hallucination, which refers to the part of descriptions that are inconsistent with the image content, such as claiming the existence of something not present in the image or describing incorrectly in terms of the counts, positions, or colors of objects in the image. Please rate the responses of the assistants on a scale of 1 to 10, where a higher score indicates better performance, according to the following criteria: 1: Accuracy: whether the response is accurate with respect to the image content. Responses with fewer hallucinations should be given higher scores. 2: Detailedness: whether the response is rich in necessary details. Note that hallucinated descriptions should not count as necessary details. Please output the scores for each criterion, containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. Following the scores, please provide an explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgement. [Assistant 1] The intended effect of this painting might be to create a humorous and whimsical twist on the famous Mona Lisa artwork by featuring a dog dressed in her attire, which can evoke feelings of amusement and curiosity. [End of Assistant 1] [Assistant 2] The intended effect of this painting is to create a humorous and whimsical scene by combining a dog with a human-like appearance, such as a Renaissance-style painting. The dog is wearing a Renaissance-style dress and a hat, which adds a touch of humor and creativity to the image. The painting might be intended to evoke amusement and curiosity in the viewer, as it is not common to see a dog dressed in such a manner. The artist might have wanted to challenge the viewer’s expectations and create a unique, memorable piece of art. [End of Assistant 2] Output format: Accuracy: <Scores of the two answers> Reason: Detailedness: <Scores of the two answers> Reason: Response: ”9”: ”**Accuracy: 8 9** *Reason:** Assistant 1 accurately describes the image as a humorous twist on the Mona Lisa featuring a dog, but lacks specific details about the attire. Assistant 2 provides a more detailed and accurate description, mentioning the Renaissance-style dress and the humorous intent, which aligns well with the image content. *Detailedness: 6 9** *Reason:** Assistant 1 gives a brief description focusing on the humorous aspect but lacks detail about the attire and style. Assistant 2 offers a richer description, noting the Renaissance-style dress and hat, and elaborates on the artistic intent, providing a more comprehensive understanding of the image.”
Table 9: Prompt for vision-based evaluation, note the image is processed by GPT-4o for evaluation.