### v3

# Through the Magnifying Glass: Adaptive Perception Magnification for Hallucination-Free VLM Decoding

Abstract
Existing vision-language models (VLMs) often suffer from visual hallucination, where the generated responses contain inaccuracies that are not grounded in the visual input. Efforts to address this issue without model finetuning primarily mitigate hallucination by contrastively reducing language biases or amplifying the weights of visual embedding during decoding. However, these approaches remain limited in their ability to capture fine-grained visual details. In this work, we propose the Perception Magnifier (PM), a novel visual decoding method that iteratively isolates relevant visual tokens based on attention and magnifies the corresponding regions, spurring the model to concentrate on fine-grained visual details during decoding. By magnifying critical regions while preserving the structural and contextual information at each decoding step, PM allows the VLM to enhance its scrutiny of the visual input, hence producing more accurate and faithful responses. Extensive experimental results demonstrate that PM not only achieves superior hallucination mitigation but also enhances language generation while preserving strong reasoning capabilities.
Introduction
The evolution of large vision-language models (VLMs) (Liu et al. 2023; Xue et al. 2024; Wang et al. 2023; Bai et al. 2023; Yang et al. 2023b) has revolutionized the integration of computer vision and natural language processing, enabling seamless interaction between visual and textual modalities. These models effectively translate complex visual inputs into coherent and contextually relevant textual descriptions, driving advancements across a wide range of applications (Wu et al. 2023; Driess et al. 2023; Wang et al. 2022). Fueled by innovations in model architectures (Desai and Johnson 2021; Eichenberg et al. 2022), training methodologies (Radford et al. 2021; Liu et al. 2023; Dai et al. 2023), and diverse multimodal datasets (Schuhmann et al. 2024; Sharma et al. 2018), large VLMs continue to expand their capabilities and adaptability, solidifying their role as foundational tools in modern AI systems.
Despite their outstanding abilities, VLMs are prone to hallucination, where generated outputs fail to align with the visual input, with inaccuracies or unfaithful content. While some hallucinations stem from VLMs adopting the same architecture as pretrained large language models (LLMs) and inheriting their intrinsic limitations (Liu et al. 2024c), the integration of visual and language features can also lead to hallucinations caused by over-reliance on language priors or limited visual representation capacity.
Figure 1: We propose Perception Magnifier (PM), a hallucination-mitigating decoding framework that locates and magnifies key visual regions based on inference attention patterns, enabling feature processing with finer details. For instance, PM locates the attention pattern on the top-right corner, and magnifies the glass bottle to it more discernible to the VLM.
Various approaches have been proposed to mitigate hallucination, ranging from training-end techniques, such as debiased dataset (Liu et al. 2024a; Zhang et al. 2025b) or enlarged visual resolution (Zhai et al. 2023; Li et al. 2024), to inference end methods such as post-processing for hallucination removal (Yin et al. 2024; Zhou et al. 2024) and specialized decoding techniques (Li et al. 2023b; Huang et al. 2024; Leng et al. 2024). Among these, hallucination-mitigation decoding methods have gained significant attention for their ability to improve visual faithfulness without requiring additional finetuning of the model. These methods primarily aim to enhance the contribution of visual information during model inference to mitigate hallucination. This enhancement can be achieved explicitly by increasing the weight of visual embeddings (Liu, Zheng, and Chen 2025; Zhu et al. 2024), prompting key regions (Yu, Yu, and Wang 2024; Yang et al. 2024), or implicitly by contrastively penalizing outputs that exhibit vague (Leng et al. 2024; Lee, Tsai, and Chiu 2024), absent (Zhang et al. 2024; Favero et al. 2024), or confused visual interpretations (Wang et al. 2024a; Chen et al. 2024a). However, these methods face limitations in scenarios where insufficient visual resolution causes object misinterpretation.
Inspired by the observation that VLMs allocate marginal attention to many visual tokens during inference and that patch-based architectures struggle with scale-sensitive visual details, we propose the Perception Magnifier (PM), a visual decoding approach that adaptively modulates spatial resolution. As shown in Fig. 1, PM magnifies critical regions for higher-resolution processing while spatially compressing rather than discarding less relevant areas, thus enhancing fine-grained detail recognition without compromising structural visual context or intrinsic reasoning ability.
Specifically, we introduce a gradient-free approach to iteratively identify high-attention visual tokens relevant to generation across deep layers and aggregate their attention scores into a token-level heatmap. This heatmap is post-processed to enhance coverage and smoothness, then upsampled to construct a pixel-level perception map. Guided by this map, PM refines the visual input by magnifying relevant regions while compressing less critical areas, implemented as a resampling process (Zheng et al. 2019; Jia et al. 2022). This dynamic adjustment allows the model to iteratively refine its focus throughout decoding, ensuring improved visual faithfulness without disrupting language reasoning.
PM is broadly applicable to interleaved VLM with unified visual sampling. Experiments on hallucination benchmarks along with GPT-4o-assisted assessments in both reference-based and vision-based settings, demonstrates that PM not only shows superior hallucination mitigation performance at question answering and open-end generation but also preserves the reasoning capabilities of VLMs. Additionally, we provide extensive ablations to analyze the contributions of each component in our method. Our contributions can be summarized as follows:
  • We propose the Perception Magnifier (PM), a decoding technique that mitigates hallucination by adaptively magnifying visual regions of interest during decoding.
  • We introduce a novel attention grounding method that iteratively aggregates attentions across layers to generate a perception map, which guides region magnification during decoding.
  • We conduct experiments with extensive evaluations, demonstrating the effectiveness of PM in mitigating hallucination while preserving reasoning capabilities.
Related Work
Hallucination in Vision-Language Models
Modern vision-language models (VLMs) typically connect vision encoders with large language models (LLMs) via either cross-attention mechanisms (Li et al. 2023a; Dai et al. 2023; Alayrac et al. 2022) or projector-based token concatenation (Liu et al. 2023; Bai et al. 2023; Koh, Salakhutdinov, and Fried 2023), with the latter dominating recent designs (Liu et al. 2024b; Hong et al. 2024; Xue et al. 2024) due to its simplicity and efficiency. While such designs align visual and textual modalities effectively, hallucination remains a persistent issue, arising not only from inherited LLM limitations (Liu et al. 2024c) but also from modality imbalance and vision encoder constraints.
To mitigate hallucination, prior efforts enhance visual grounding by prompting input images (Yu, Yu, and Wang 2024; Yang et al. 2024; Gizdov, Ullman, and Harari 2024), boosting visual token influence (Zhang et al. 2024; Zhu et al. 2024; Liu, Zheng, and Chen 2025), or improving vision module capacity (Zhai et al. 2023; Liu et al. 2024b; Chen et al. 2024b). Data-centric strategies include debiasing (Kim et al. 2025; Liu et al. 2024a; Zhang et al. 2025b; Ouali et al. 2025; Yue, Zhang, and Jin 2024), contrastive training (Jiang et al. 2024), and evaluation techniques (Kaul et al. 2024; Guan et al. 2024; Wang et al. 2024b; Li et al. 2023c; Lovenia et al. 2024; Jing et al. 2024), along with post-hoc filtering (Yin et al. 2024; Zhou et al. 2024). While these approaches tackle hallucination from multiple angles, few methods focus on enhancing visual recognition capabilities beyond increasing resolution. We propose a decoding strategy that magnifies semantically important regions without disrupting spatial coherence, improving perception during generation.
Hallucination Mitigation Inference Strategies
Some decoding-end approaches reduce hallucination by modifying next-token generation. Contrastive decoding (CD) (Li et al. 2023b) amplifies expert model logits while suppressing outputs from a biased source, typically a smaller or noisier model. Chuang et al. streamlines this by using early-exit logits from the expert model itself (2024). OPERA (Huang et al. 2024) penalizes beam outputs that heavily attend to generic aggregation tokens. VCD (Leng et al. 2024) adapts CD for VLMs by defining biased outputs from noisy image inputs. Later works extend this by deriving biased outputs by removing image inputs (Zhang et al. 2024; Favero et al. 2024) or enhance focus on textual feature (Wan et al. 2025), masking key visual tokens (Huo et al. 2024), or synthesizing biased outputs through finetuning (Chen et al. 2024a) or prompting (Wang et al. 2024a). Others enhance expert logits by upweighting visual tokens (Liu, Zheng, and Chen 2025; Zhu et al. 2024) or selecting focal views via external detectors (Chen et al. 2024c). Lee, Tsai, and Chiu fuse multiple biased outputs using weighted voting (2024). While these decoding strategies mitigate hallucination by penalizing visually inconsistent outputs, over-suppressing biased signals may hinder coherence and reasoning. Alternatively, some inference-time methods enhance visual grounding by applying masks or bounding boxes (Yu, Yu, and Wang 2024; Yang et al. 2024, 2023a), but still struggle to recognize fine-grained details. ViCrop (Zhang et al. 2025a) improves upon this by cropping and resizing key visual regions as auxiliary inputs. However, it disrupts spatial structure and may impair tasks requiring holistic understanding, and doubles visual tokens with more compute. In contrast, our method magnifies semantically important regions while compressing less relevant areas, preserving spatial integrity and supporting both visual grounding and coherent generation.
Figure 2: Structural diagram of the proposed PM method. At each decoding step, we aggregate multi-layer attentions into a token-level heatmap $\mathcal{H}^{(i)}$. By iteratively masking the most attended tokens and re-forwarding, we collect heatmaps $\{\mathcal{H}^{(i)}\}$ and combine them into $\mathcal{H^{*}}$ once the process ends, and post-processing $\mathcal{H^{*}}$ to obtain $\mathcal{P}$ (yellow block). The magnifier resamples the visual input according to $\mathcal{P}$, producing $\hat{I}$ (purple block). The VLM can then generate the next token with enhanced focus on critical regions with finer details.
Methods
In this paper, we propose the Perception Magnifier (PM), a novel approach that iteratively identifies and enhances key regions of the input image to refine visual input for less hallucinative token generation in VLMs. As illustrated in Fig. 2, PM iteratively identifies highly attended visual tokens, blocks them, and re-forwards the sequence until no tokens receive significant attention. The identified tokens are then aggregated based on their attention scores to construct a perception map $\mathcal{P}$. This map is used to magnify the original image $I$ via a sampling process, $\hat{I}^{i}=\mathcal{T}(I,\mathcal{P})$. The refined image $\hat{I}^{i}$ then serves as the refined visual input for generating the next-step output logit, $O^{i+1}=\text{VLM}(\hat{I}^{i},O^{i})$, where $O^{i}$ denotes the probability logits produced at step $i$. The following sections detail the perception map construction, the iterative refinement process for improving the coverage of perception map, and the attention-based magnification processed designed in the proposed PM method.
Perception Map Construction
In this section, we first introduce a simple method to construct a basic perception map to guide the magnifier. This method applies to mainstream interleaved VLMs, where patchified visual embeddings are integrated with textual tokens into a unified input sequence processed by the VLM.
The transformer processes the input sequence via self-attention. For a single attention head, the attention weight from token $x_{i}$ to $x_{j}$ is computed as $\mathrm{Attn}_{i,:}=\text{Softmax}\left(Q_{i}K_{j}^{\top}/\sqrt{d_{k}}\right)$, where $Q_{i}$ and $K_{j}$ are the query and key vectors of tokens $x_{i}$ and $x_{j}$, respectively, and $d_{k}$ is the key dimension. The corresponding value $V_{j}$ is then weighted by $\mathrm{Attn}_{i,j}$ to obtain the hidden representation of $x_{i}$: $z_{i}=\sum_{j}\mathrm{Attn}_{i,j}\,V_{j}$. In practice, transformers stack multiple layers and heads, producing attention weights across different levels. These attention patterns can be represented as a multi-dimensional tensor: $\mathrm{Attn}\in\mathbb{R}^{N_{l},N_{h},H,W}$, where $N_{l}$ and $N_{h}$ denote the number of layers and attention heads, respectively, and $H\times W$ represents the spatial dimensions of visual tokens, which can be reshaped into a 2D feature map. PM utilizes the $\mathrm{Attn}$ cached during forward to construct a token-level heatmap, highlighting the most relevant visual regions based on the model attention.
Empirically, we find that attention maps from middle layers more accurately localize relevant objects than those from final layers. This aligns with recent findings showing that deep networks tend to aggregate and process visual information in early layers (Darcet et al. 2024), and that intermediate features can already support strong performance (Chuang et al. 2024). Hence, we construct the perception map by aggregating self-attention matrices from middle-to-late layers $l\geq\mathcal{L}$ and applying a max pooling across all heads $h$ to highlight the impact of more visually important regions. Formally, the token-level heatmap $\mathcal{H}$ can be obtained by:
$\mathcal{H}=\sum_{l=\mathcal{L}}^{N_{l}}\max_{h\in{1,\ldots,N_{h}}}\mathrm{Attn}_{l,h}.$ (1)
Such $\mathcal{H}\in\mathbb{R}^{H,W}$ reflect the intermediate reasoning of VLM and highlight the salient regions identified.
Post-Processing to Pixel Space.
The basic perception map constructed above without iterative refinement could already be upsampled to guide the magnifier. As post-processing before up-sampling the token-level heatmap $\mathcal{H}$ to pixel-level perception map, we first refine the heatmap with methods similar to (Yu, Yu, and Wang 2024) to obtain a more smooth and well-distributed attention distribution for better magnification quality. Firstly, we normalize $\mathcal{H}$ by scaling it to unit interval, obtaining $\mathcal{H}^{\prime}$. Then, we amplify the variance of the normalized map by a scaling coefficient of $\alpha$, and then compress its values back into the range $[0,1]$ using a sigmoid function, resulting in $\mathcal{H}_{\text{proc}}$:
$\mathcal{H}_{\text{proc}}=\text{sigmoid}\!\left(\alpha\cdot\,\frac{\mathcal{H}^{\prime}-\mu(\mathcal{H}^{\prime})}{\sigma(\mathcal{H}^{\prime})}\right).$ (2)
Finally, we apply a uniform smoothing filter with a kernel size of $k$ to the enhanced heatmap, followed by bilinear upsampling to obtain the pixel-level perception map $\mathcal{P}$. As shown in Fig. 2, this post-processing step refines the heatmap by amplifying low-magnitude attention values, ensuring that small but potentially important regions are also considered for magnification.
Iterative Refinement
While the basic perception map already provides a simple but effective guide for magnification, it often misses some key objects’ regions when they appear multiple times. This stems from how deep vision models compress fine-grained features into a small set of tokens, commonly referred to as information registers (Darcet et al. 2024; Reddy 2024). VLMs concentrate attention to these tokens and reduce sensitivity to semantically similar but spatially distinct regions.
To improve coverage of the perception map, we introduce an iterative refinement mechanism that progressively reveals overlooked areas by suppressing highly attended tokens from prior passes. Nonetheless, directly masking key visual regions at the image level would require fine-grained localization and re-tokenization of visual tokens, making it computationally expensive. Instead, we perform attention blocking at the token level (Alg. 1, lines 4–5), which still blocks aggregated tokens, efficiently mitigates above challenges, and integrates easily into existing VLM pipelines.
At each iteration, we forward the input with an attention mask $M^{(i)}$, initialized as a vector of ones (line 1), and extract the token-level heatmap $\mathcal{H}^{(i)}$ (line 5). To identify visually dominant tokens $T^{(i)}$, we apply a simple clustering algorithm that partitions attention values into two groups (line 7), and select tokens in the one with the higher centroid as $T^{(0)}$. This leverages the observation that relevant tokens typically receive significantly higher attention than others. We then update the mask to exclude these tokens in the next iteration (line 11), encouraging the model to reallocate attention to previously under-attended regions.
The process continues until the total attention in the new heatmap drops below a threshold $\beta$ or a maximum number of iterations is reached (lines 8–10), indicating that no additional high-attention regions remain. We aggregate attention maps across all iterations and normalize them to obtain the final refined heatmap $\mathcal{H}^{*}$ (lines 13), which is then post-processed to produce the pixel-level perception map $\mathcal{P}$ (line 14). This refined map captures a broader and more accurate set of visually relevant regions, offering better guidance for magnification-enhanced decoding in downstream tasks.
Algorithm 1 Iterative Refinement on the Perception Map
1:Tokenized input $(I,Q)$, Visual Language Model (VLM), threshold $\beta$, max iterations $T_{\max}$
2:Initialize attention mask $M^{(0)}\leftarrow\mathbf{1}$
3:Initialize heatmap sum $\mathcal{H}^{*}\leftarrow\mathbf{0}$, count mask $\mathcal{C}\leftarrow\mathbf{0}$
4:for $i=0$ to $T_{\max}$ do
5:  $O^{(i)},\text{Attn}^{(i)}\leftarrow\text{VLM}(I,Q,M^{(i)})$
6:  Compute heatmap: $\mathcal{H}^{(i)}=\sum_{l\in\mathcal{L}}\max_{h}\mathrm{Attn}^{(i)}_{l,h}$
7:  Update $\mathcal{H}^{*}\leftarrow\mathcal{H}^{*}+\mathcal{H}^{(i)}$, $\mathcal{C}\leftarrow\mathcal{C}+\mathbf{1}_{M^{(i)}}$
8:  Identify $T^{(i)}\leftarrow\text{Cluster}(\mathcal{H}^{(i)})[0]$
9:  if $\sum\mathcal{H}^{(i)}<\beta$ then
10:   break
11:  end if
12:  Mask out $T^{(i)}$: $M^{(i+1)}\leftarrow M^{(i)}\cap\neg T^{(i)}$
13:end for
14:Normalize heatmap: $\mathcal{H}^{*}\leftarrow\mathcal{H}^{*}\div\mathcal{C}$
15:Post-process $\mathcal{H}^{*}$ to obtain pixel-level map $\mathcal{P}$
16:return Refined perception map $\mathcal{P}$
Attention-Based Magnification
Inspired by (Zheng et al. 2019), we develop the perception magnifier as an attention-based sampling process. The perception map $\mathcal{P}$ is treated as a mass function, where regions with higher perception values are more likely to be sampled. Specifically, we first decompose $\mathcal{P}$ into marginal distributions along the horizontal and vertical dimensions and compute their cumulative distributions $\mathcal{F}_{x}(n)$ and $\mathcal{F}_{y}(n)$ as:
$\begin{split}\mathcal{F}_{x}(n):=\sum_{j=1}^{n}\max_{1\leq i\leq w}\mathcal{P}_{i,j}\quad\mathcal{F}_{y}(n):=\sum_{i=1}^{n}\max_{1\leq j\leq h}\mathcal{P}_{i,j},\end{split}$ (3)
where $w$ and $h$ denote the width and height of the perception map, respectively. Next, we apply inverse transform sampling to remap the original image coordinates to resampled coordinates, ensuring structure-preserving magnification along both dimensions. The magnified image is then constructed as:
$\hat{I}=\mathcal{T}(I,\mathcal{P})_{i,j}=\text{Interp}\big{(}I,\mathcal{F}_{x}^{-1}(i),\mathcal{F}_{y}^{-1}(j)\big{)},$ (4)
where $\mathcal{F}^{-1}(\cdot)$ denotes the inverse function of $\mathcal{F}(\cdot)$, which remaps uniformly spaced output coordinates to a perception-aware warped space. To ensure smooth spatial mapping, we use bilinear interpolation as the Interp method in the equation, balancing efficiency and effectiveness, though other interpolation methods could also be applied. The magnified image $\hat{I}$ replaces the original visual input $I$ and is forwarded through the VLM with the same textual input to obtain updated model output logits $\hat{O}$. While perception-magnified logits $\hat{O}$ can be proportionally combined or contrastively compared with the original model output $O$ for next-token generation, we directly sample from $\hat{O}$ after the softmax function, as it already provides an enhanced representation of the visual content. The entire perception magnification process is applied at each auto-regressive decoding step, enabling the model to adaptively focus on different regions of the visual input while preserving structural integrity, hence enhancing perception ability and mitigating hallucination.
Object-level Attribute-level MME MME POPE Accuracy
Existence Count Position Color Percep* Cognition AOKVQA GQA COCO AVG*
Greedy 195.00 143.33 128.33 163.33 630.00 357.86 84.10 84.39 85.29 84.59
OPERA 195.00 143.33 128.33 163.33 630.00 310.36 84.09 84.43 85.29 84.60
VCD 190.00 143.33 120.00 155.00 608.33 327.50 84.47 83.13 85.61 84.40
ICD 190.00 148.33 123.33 153.33 615.00 348.21 84.23 84.89 84.92 84.68
SID 195.00 155.00 126.67 160.00 636.67 327.86 84.39 84.17 85.81 84.79
VDD 190.00 140.00 133.33 165.00 628.33 322.64 86.73 85.52 86.71 86.32
M3ID 190.00 150.00 133.33 166.67 640.00 339.36 86.52 85.46 86.73 86.24
PAI 190.00 155.00 133.33 170.00 648.33 353.93 85.32 83.43 86.66 85.14
IBD 190.00 160.00 133.33 170.00 653.33 352.14 86.37 84.80 86.65 85.94
API-C 190.00 150.00 128.33 165.00 633.33 317.50 86.87 85.04 87.31 86.41
API-L 190.00 150.00 133.33 168.33 641.67 325.00 86.52 84.30 86.60 85.81
ViCrop-R 190.00 163.33 105.00 175.00 633.33 335.00 85.53 85.19 87.43 86.05
ViCrop-G 195.00 153.33 96.67 175.00 620.00 350.29 86.09 85.18 87.31 86.19
[…]
Table 1: Evaluation results for MME Perception Scores (left), Cognition Scores (mid), and POPE Accuracy (right). Columns marked with * indicate the overall hallucination scores for the benchmark. The result for BEAM is omitted for its equivalence with OPERA on single-token QA. API and ViCrop each propose multiple variants, denoted by postfixes (e.g., -C or -L), with details provided in the Supplementary. The proposed PM method demonstrates superior performance on both benchmarks.
Experiments and Results
Datasets and Baselines.
We evaluate our method on three benchmarks: MME (Fu et al. 2023), POPE (Li et al. 2023c), and LLaVA-Bench (Liu et al. 2023), all commonly used for assessing multimodal hallucination. For comparison, we include various decoding or inference strategies, including greedy decoding, VCD (Leng et al. 2024), VDD (Zhang et al. 2024), M3ID (Favero et al. 2024), IBD (Zhu et al. 2024), PAI (Liu, Zheng, and Chen 2025), ICD (Wang et al. 2024a), SID (Huo et al. 2024), beam search (Sutskever, Vinyals, and Le 2014), OPERA (Huang et al. 2024), API (Yu, Yu, and Wang 2024), and ViCrop (Zhang et al. 2025a). Further details are provided in the Supplementary.
Implementation Details.
We conduct our experiments using LLaVA-1.5 7B (Liu et al. 2024b). We start aggregating attentions from layer $\mathcal{L}=12$, set the scaling coefficient and kernel size to $\alpha=10$, $k=3$, and set the attention threshold for iterative refinement to $\beta=0.3$.
Quantitative Analysis
Table 1 reports results on MME and POPE. POPE centers on existence-based queries, where most methods benefit from reinforcing visual token importance. MME is more challenging, with its perception subset targeting hallucination and its cognition subset requiring higher-level reasoning. While our method focuses on mitigating hallucination, we report cognition results for completeness, noting that most baselines underperform on this subset as modifications on the embedding and logit could harm the VLM reasoning.
Among baselines, CD-based methods suppress hallucinated logits but often fail when visual detail remains ambiguous across both logits. Embedding-level methods such as PAI and IBD improve perception by reinforcing visual features across multiple layers. However, they may still struggle when relevant regions are too small or diffuse to be effectively represented in the ViT backbone. API suppresses irrelevant regions via masking, but does not enhance salient visual content. ViCrop isolates key areas via cropping, which may hinder spatial understanding due to loss of context and potential ambiguity from dual-image inputs, resulting in limited MME gains. In contrast, our method magnifies the relevant regions while preserving surrounding context and internal states of the VLM. This preserves the embedding distribution and enhances fine-grained recognition without compromising spatial coherence or reasoning, leading to state-of-the-art results across both benchmarks.
In addition, all methods exhibit varying degrees of degradation on the MME cognition subset, whereas ours does not. A detailed analysis of how and why different methods affect the cognitive ability is provided in the Supplementary.
Question:
Is there a chair
in the image?
Greedy: No
PM: Yes
Question:
Is there a tree
in the image?
Greedy: Yes
PM: No
Figure 3: Qualitative examples on the POPE benchmark showing the input question, the original image, the raw and processed perception map $\mathcal{H}^{*}$ and $\mathcal{P}$, and the perception-magnified image, arranged from left to right. The answers generated by greedy decoding and PM are compared, with correct answers highlighted in green and incorrect ones in red.
GPT-4o Assisted Evaluation
LLaVA-Bench COCO In-the-Wild
Ref $\uparrow$ Acc $\uparrow$ Det $\uparrow$ Ref $\uparrow$ Acc $\uparrow$ Det $\uparrow$
Greedy 81.15 92.69 95.76 56.94 71.83 79.95
BEAM 83.17 94.75 91.60 58.94 74.71 78.25
OPERA 82.99 96.12 93.61 58.44 71.81 77.92
VCD 82.00 92.21 94.39 57.17 69.4 81.22
ICD 83.29 93.91 95.46 54.43 66.89 74.81
SID 81.10 97.35 95.80 59.49 70.93 81.63
VDD 81.38 95.62 102.75 61.60 78.22 83.01
M3ID 83.02 96.41 95.78 61.88 77.49 82.83
PAI 79.04 92.06 90.59 52.29 69.74 72.49
IBD 85.29 95.73 95.45 61.59 77.91 82.86
[…]
Table 2: Results on LLaVA-Bench (%). Ref denotes reference aggrement, while Acc and Det represent accuracy and detailedness in vision-based evaluation.
Previous studies (Huang et al. 2024; Zhu et al. 2024) have demonstrated the reliability of GPT-4-assisted evaluation for hallucination assessment. We extend their evaluation protocol and experiment on LLaVA-Bench to benchmark the open-end generation quality. Overall, GPT-4o scores responses based on reference agreement (Ref), visual accuracy (Acc), and descriptive detail (Det), depending on the evaluation setting. We omit API and ViCrop as they are not applicable to open-ended generation. Details on the GPT-4o metrics and prompts can be found in the Supplementary.
As shown in Tab. 2, the proposed PM achieves the most competitive overall performance on open-end generation tasks. While its detailedness on COCO and accuracy on the In-the-Wild split are slightly lower than a few existing method, PM still demonstrates advantage in hallucination mitigation, particularly in reference-based evaluation. These results indicate that even in open-end decoding where the model does not always need to focus on specific regions at each step, the proposed adaptive perception magnifier substantially enhances output quality, demonstrating its effectiveness in reducing hallucinations.
Qualitative Analysis
Figure 3 presents qualitative examples illustrating the input question, the original image, the raw and processed perception map $\mathcal{H}^{*}$ and $\mathcal{P}$, and the perception-magnified image. The results demonstrate that the proposed method effectively enlarges small objects relevant to the question while preserving fine details. In the first example, the VLM initially attends to the top-left region, recognizing the potential presence of a chair but failing to identify it. With the magnified input, the chair is presented at a higher spatial resolution, enabling the VLM to correctly recognize it. Similarly, in another case, the VLM misidentifies an object behind a door window as a tree. However, the perception-magnified input allows the model to re-examine the object with greater detail, preventing hallucinated predictions.
Overall, the qualitative results suggest that the proposed method enhances the perceptibility of fine-grained visual details in regions critical for next-token generation. By enabling the model to examine these areas at a higher resolution, our approach effectively reduces hallucinations and improves overall model performance.
Ablations
MME-Perception Object-level Attribute-level Percep
Existence count position color Total*
Greedy 195.00 143.33 128.33 163.33 630.00
PM w/o IR & MLA 195.00 155.00 120.00 170.00 640.00
PM w/o MLA 195.00 155.00 125.00 170.00 645.00
PM w/o IR 190.00 175.00 135.00 165.00 665.00
[…]
Table 3: Ablations on iterative refinement (IR) and multi-layer aggregation (MLA). Both contribute to performance improvements compared to their respective baselines.
Iterative Refinement.
We ablate the effect of iterative refinement in constructing the perception map. As shown in Tab. 3, the refined maps consistently outperform their non-iterative counterparts, confirming the benefit of this strategy in improving visual localization. Iterative refinement helps address tokens overlooked from the early-stage aggregation of visual features, resulting in more accurate coverage of relevant regions. Although some residual leakage from masked tokens remains due to information exchange in the ViT (Dosovitskiy et al. 2021) visual feature extractor, our approach still achieves decent grounding of key areas with minimal computational overhead, making it an effective and efficient solution for guiding the magnifier.
MME-Perception Object-level Attribute-level Percep
Existence Count Position Color Total*
(a) Perception Map Construction Methods
API CLIP 190.00 155.00 131.67 170.00 646.67
API LLaVA 190.00 165.00 126.67 170.00 651.67
ViCrop-R 195.00 163.33 116.67 170.00 645.00
ViCrop-G 195.00 158.33 116.67 170.00 640.00
Rel Map 195.00 175.00 130.00 155.00 655.00
[…]
(b) Perception-Guided Magnification Methods
Blurring 185.00 158.33 111.67 175.00 630.00
Bounding Box 180.00 160.00 130.00 170.00 640.00
Masking 195.00 155.00 128.33 170.00 648.33
ViCrop 195.00 153.33 123.33 175.00 646.67
[…]
Table 4: Ablation results across two groups: (a) comparing different methods for constructing perception maps; (b) comparing prompting techniques that highlight key regions based on the map. IR denotes iterative refinement.
Multi-Layer Aggregation.
We ablate the effect of aggregating attention across multiple layers when constructing the perception map. As shown in Tab. 3, multi-layer aggregation yields significantly better coverage of relevant visual regions compared to using attention from a single layer.
While middle-to-deep layers generally focus more on visual content, the most informative attention varies across layers and samples. In many cases, only a subset of intermediate layers exhibits structured visual attention, while others focus primarily on sink tokens (Xiao et al. 2024). Aggregating across layers reduces this variability, ensuring that key visual cues are reliably captured for guiding the magnifier.
Comparison with Different Maps.
Attention-based visualization methods (Chefer, Gur, and Wolf 2021; Stan et al. 2024) have been proposed to improve VLM interpretability by combining attention weights with gradients. We adapt several representative techniques—including CLIP-based Rel(evance) Maps, API (Yu, Yu, and Wang 2024), and ViCrop (Zhang et al. 2025a), to construct a map to guide the magnifier, and compare them to our approach w/o IR using identical post-processing. As shown in Tab. 4, our method achieves superior coverage of relevant visual regions. Attribution maps based on CLIP often reflect global similarity rather than token-level saliency, while API and ViCrop variants aggregate attention from a single layer, resulting in lower stability and diluted focus. Backpropagation-based methods (e.g., Rel Map, ViCrop-G) offer fine-grained attribution but contribute little to magnification quality and are computationally expensive. In contrast, the proposed strategy, even w/o the IR process, preserves critical visual cues with better spatial consistency and minimal overhead, making it well-suited for efficient perception-guided decoding.
Image Prompting Methods.
Various methods have been proposed to “prompt” images to emphasize specific regions (Yang et al. 2024; Zhang et al. 2025a), including add a bounding box, apply blurring effect, masking out the remaining regions, or crop and resize the key area as additional input (ViCrop). We benchmark these methods as alternatives to our magnifier, evaluating their effectiveness in highlighting key visual regions when combined with the constructed perception map. As shown in Tab. 4, our magnification approach more effectively leverages the perception map to enhance the perception ability of the VLM.
This difference arises because magnification not only highlights important regions but also reveals additional fine-grained details within limited resolution, allowing the model to capture subtle visual patterns. In contrast, many prompting methods merely emphasize region importance without increasing effective resolution, limiting their impact on fine-grained recognition. Meanwhile, ViCrop also yields certain improvement, as it also enhances local details with resized key regions. However, ViCrop discards contextual information and confuses the VLM by supplying the additional resized image, leading to reduced effectiveness in tasks such as position-related and counting questions.
[…]
Figure 4: Hyperparameter analysis of $\alpha$ and $\beta$ on MME. PM demonstrates robust performance across different settings.
Hyperparameter Analysis on $\alpha$ and $\beta$.
We analyze the effect of $\alpha$ and $\beta$, which control magnification strength and the stopping criterion for iterative refinement, respectively. Increasing $\alpha$ enhances heatmap contrast and strengthens magnification, while decreasing $\beta$ increases refinement steps, improving coverage but at higher computational cost (capped at 8 iterations). As shown in Fig. 4, smaller $\alpha$ pairs well with low $\beta$, which already ensures broad region coverage, reducing the need for strong magnification. Overall, the perception magnifier remains robust across a wide range of hyperparameters, consistently identifying key visual regions. Notably, using a larger $\beta$ approximates the performance of the basic PM w/o iterative refinement, yet substantially mitigates hallucinations with only $2\times$ the compute.
Conclusion
We introduce the Perception Magnifier (PM), a visual decoding method that refines image inputs to mitigate hallucinations in vision-language models. PM adaptively enlarges visually critical regions identified by multi-layer attention iteratively, providing enhanced effective resolution on relevant finer-grained details at each decoding step. Experiments demonstrate that PM significantly reduces hallucinations at VQA and open-end generation, while preserving reasoning capabilities. Future work may focus on optimizing the efficiency for PM at long-text generation, and mitigate shape distortions in the magnification process.
Supplementary Materials
Method AOKVQA GQA COCO
Adversarial Popular Random Adversarial Popular Random Adversarial Popular Random
Greedy 78.67 84.93 88.70 80.40 83.27 89.50 83.50 85.57 86.80
Opera 78.60 84.93 88.73 80.43 83.33 89.53 83.47 85.57 86.83
VCD 79.10 84.93 89.37 79.63 81.70 88.07 83.53 85.63 87.67
ICD 79.07 85.40 88.23 80.97 84.77 88.93 83.20 85.23 86.33
SID 78.90 85.40 88.87 80.60 83.33 88.57 83.90 86.07 87.47
VDD 81.03 88.27 90.90 82.00 85.27 89.30 85.00 87.00 88.13
M3ID 80.87 87.97 90.73 81.67 84.93 89.77 85.10 86.93 88.17
PAI 79.63 86.27 90.07 79.90 82.30 88.10 85.03 86.97 87.97
IBD 80.60 87.47 91.03 81.30 84.10 89.00 84.90 86.97 88.07
API-C 81.00 88.20 91.40 81.33 84.27 89.53 85.47 87.53 88.93
API-L 80.80 87.70 91.07 81.00 83.57 88.33 84.53 86.97 88.30
ViCrop-R 80.00 86.37 90.23 81.60 83.83 90.13 85.60 87.70 89.00
ViCrop-G 80.97 86.77 90.53 81.67 83.93 89.93 85.13 87.70 89.10
[…]
Table 5: Evaluation results for POPE accuracy with LLaVA 1.5.
Datasets
We carry out our experiments on three benchmarks popular for examining the perception ability and hallucination of VLMs.
MME (Fu et al. 2023) benchmark is designed to comprehensively assess both perception and cognition abilities of VLMs across 14 distinct subtasks on 2,370 questions. In our study, we focus primarily on the Perception subset of MME following (Leng et al. 2024) to evaluate the effectiveness of our method in controlling hallucinations.
POPE (Polling-based Object Probing Evaluation) (Li et al. 2023c) is a 9 $\times$ 3,000 question benchmark for evaluating object hallucination in VLMs. It formulates hallucination evaluation as a binary classification task by posing yes/no questions about object presence, addressing limitations of CHAIR (Rohrbach et al. 2018), which relies on exact text matching, and providing a more robust evaluation less sensitive to prompt variations.
LlaVA-Bench (Liu et al. 2023) is a benchmark designed to evaluate VLMs in real-world visual chat scenarios. It comprises two components: LLaVA-Bench (COCO), utilizing images from the COCO dataset, and LLaVA-Bench (In-the-Wild), a more challenging subset featuring 24 diverse images—such as indoor and outdoor scenes, memes, paintings, and sketches—paired with 60 detailed, manually curated questions. The benchmark covers various vision-language tasks, making it suitable for assessing model generalization. In our study, we employ both COCO and In-the-Wild splits to comprehensively evaluate different methods on hallucination control, cognitive abilities, and performance in open-ended question answering.
Baselines
This section summarizes the high-level idea of methods used as comparisons in the experiments. We apply these methods on our benchmarks with the default hyperparameters provided in the original paper.
Beam Search explores multiple output candidates by iteratively expanding the most probable sequences and selecting the top-$k$ hypotheses at each step, resulting in improved answer diversity but at the cost of increased computational complexity proportional to the beam size. We evaluated all beam-based methods with a beam size of 4.
OPERA (Huang et al. 2024) is a beam-search-based method, examining the presence of a hallucinate beam and pruning the beam based on its attention on information aggregation patterns.
VCD (Leng et al. 2024) introduces noise into the input image and processes it through the VLM to generate a hallucinative output. By leveraging contrastive decoding (CD) to penalize high-probability predictions that align with the hallucinative output, VCD reduces the likelihood of hallucinations in the final prediction, resulting in more reliable responses.
ICD (Wang et al. 2024a) generates a hallucinative output for CD-based decoding by appending “You are a confused object detector” to the input prompt.
SID (Huo et al. 2024) induces a hallucinative output for CD-based decoding by identifying key visual tokens in the image and pruning them through forward.
VDD (Zhang et al. 2024) mitigates language-induced hallucinations in VLMs by generating a hallucinative output through the removal or replacement of visual tokens with <unk> tokens.
M3ID (Favero et al. 2024) is similar to VDD but with weighted strength of incorporating contrastive corrections based on the text-image mutual information.
PAI (Liu, Zheng, and Chen 2025) enhances visual faithfulness by increasing the weights of visual tokens through attention mechanisms while penalizing outputs generated without visual inputs using CD.
IBD (Zhu et al. 2024) also biases the output toward visual inputs by adding visual attentions and refines it against the original output using CD. Additionally, it adaptively adjusts the contrastive update strength at each decoding step based on contextual needs.
API (Yu, Yu, and Wang 2024) highlights key visual regions by overlaying a gray-level mask on the input image. The mask is guided by a heatmap derived from image-text associations using a CLIP model (-C) or a LLaVA model (-L).
ViCrop (Zhang et al. 2025a) estimates a bounding box around the key region in the image and generates a new input by cropping and resizing this region. Three approaches are proposed for bounding box estimation with different level of reliance on gradient information: relative attention (-R), gradient attention (-G), and pure gradient (-P). The pure-gradient approach is not benchmarked due to its higher GPU memory consumption, which necessitates lower model precision and results in degraded performance, making it unsuitable for a fair comparison.
Computing Infrastructure
All experiments were conducted on a single machine equipped with an NVIDIA GeForce RTX 3090 GPU and an Intel(R) Core(TM) i7-10700 CPU. The system runs Ubuntu 22.04 LTS and uses CUDA version 12.1 for GPU acceleration. Experiments were implemented in Python using PyTorch (v2.4.0) and Transformers (v4.37.2).
Maintaining Cognition Ability
MME-Cognition CS NR Trans CR Total*
Greedy 112.86 70.00 107.50 67.50 357.86
OPERA 112.86 47.50 90.00 60.00 310.36
VCD 120.00 37.50 97.50 72.50 327.50
ICD 120.71 70.00 92.50 65.00 348.21
SID 122.86 55.00 85.00 65.00 327.86
VDD 110.14 58.00 68.00 86.00 322.64
M3ID 110.86 61.00 92.50 75.00 339.36
PAI 116.43 70.00 100.00 67.50 353.93
IBD 117.14 70.00 107.50 57.50 352.14
API-C 115.00 47.50 107.50 47.50 317.50
API-L 120.00 45.00 115.00 45.00 325.00
ViCrop-R 115.00 50.00 105.00 65.00 335.00
ViCrop-G 109.29 50.00 105.00 86.00 350.29
[…]
Table 6: Comparison of different hallucination control decoding methods on the MME-Cognition subset, including Commonsense Reasoning (CS), Numerical Reasoning (NR), Translation (TR), and Code Reasoning (CR).
Many hallucination control decoding methods enhance perception by increasing the contribution of visual information during token generation. While these approaches generally improve perceptual ability, they often disrupt the calibrated balance between the vision and language modalities, potentially compromising the inherent cognitive capabilities learned by the VLM during large-scale pretraining.
To investigate this, we evaluate our method and baselines on the Cognition split of MME, where answering requires reasoning beyond direct visual cues. As shown in Tab. 6, most CD-based methods that explicitly modify output logits exhibit a notable decline in cognition ability, aligning with expectations. Surprisingly, PAI and IBD, which enhance the attention weight of visual tokens through intermediate embedding modifications, do not incur a substantial loss in cognition performance. This suggests that the bias introduced by their layer-wise manipulation is relatively mild and can be compensated for by subsequent model layers. Furthermore, despite API-based methods modify the image without altering the VLM’s internal reasoning, they still lead to degraded performance. We hypothesize that the gray masking applied to images in the cognition set where backgrounds are typically uniform may introduce additional perceptual burden, ultimately impacting overall performance.
In contrast, our proposed PM method does not negatively affect cognition. Although our approach is designed solely for hallucination control rather than cognition enhancement, the total cognition score remains comparable to greedy decoding and outperforms other methods, demonstrating its robustness in preserving reasoning capabilities.
Quantitative Results on POPE
Table 5 presents the detailed evaluation results across each split of the POPE benchmark. Overall, the proposed PM consistently outperforms existing methods on all splits, demonstrating its effectiveness in mitigating hallucination in yes/no existence-based VQA. While the improvement on the ‘popular’ split of GQA is relatively smaller, this may be attributed to the inherently dense scenes and lower object saliency in the GQA benchmark, which can pose challenges for our magnification-based visual enhancement. Nevertheless, PM still achieves strong performance on this split, highlighting its robustness even under complex visual conditions. These results confirm that PM generalizes effectively across diverse data sources and multiple types of hallucination-oriented VQA.
Object-level Attribute-level MME Score POPE Accuracy
Existence Count Position Color Total* AOKVQA GQA COCO AVG*
Greedy 190.00 160.00 71.67 155.00 576.67 81.87 81.27 83.46 82.19
VCD 185.00 163.33 65.00 148.33 561.67 81.86 81.14 83.31 82.10
ICD 195.00 165.00 65.00 145.00 575.00 81.93 81.05 83.78 82.25
SID 195.00 155.00 73.33 145.00 568.33 81.57 81.40 83.43 82.13
VDD 195.00 155.00 73.33 155.00 578.33 81.99 81.74 83.03 82.26
M3ID 195.00 160.00 75.00 155.00 585.00 81.92 81.54 83.43 82.30
API-C 185.00 165.00 76.67 165.00 591.67 82.21 81.28 83.47 82.32
API-L 190.00 165.00 73.33 160.00 588.33 81.84 81.24 83.33 82.14
[…]
Table 7: Evaluation results for MME Perception Scores (left) and POPE Accuracy (right) on the Shikra VLM. Columns marked with * indicate the overall scores for the benchmark. Some methods are omitted due to limitations on computational resources and implementations. The proposed PM method consistently demonstrates superior performance on both benchmarks.
Performance on Shikra
To assess the generalizability of PM beyond LLaVA, we apply it, along with several comparative methods, to Shikra-7B (Chen et al. 2023), a VLM designed for referential dialogue. The performance is evaluated on MME and POPE using Shikra with iterative refinement threshold $\beta=1.5$. The results are presented in Tab. 7. Note that certain comparative methods are excluded either due to the lack of available implementations for Shikra or constraints on the computational resources available.
As shown in Tab. 7, the proposed method continues to improve performance on Shikra compared to the baselines, though the improvement is less pronounced than on LLaVA. This discrepancy may stem from Shikra’s lower visual resolution and fewer visual tokens, which impact the effectiveness of perception magnification. With fewer visual tokens, cases where an object is partially recognized at intermediate layers but not reflected in the final prediction become less frequent, as each token carries greater attention weight in the inference process. Therefore, the proposed method has fewer opportunities to enhance recognition. Additionally, the reduced number of visual tokens lowers the granularity of the token-level perception map, leading to less precise perception magnification, which may further limit the performance gains of PM.
This observation suggests that the proposed method could be particularly beneficial for high-resolution VLMs, where small local features are more likely to be detected at intermediate layers but fail to significantly influence the final prediction. However, verifying this hypothesis is constrained by computational resource limitations.
Overall, the results on Shikra provide strong evidence that the proposed PM method generalizes well beyond LLaVA and effectively enhances hallucination mitigation in applicable VLMs.
Question:
Is there a building
in the image?
Greedy: No
PM: Yes
Question:
Is there a bottle
in the image?
Greedy: No
PM: Yes
Question:
Is there a paper
in the image?
Greedy: Yes
PM: No
Question:
Is there a cabinet
in the image?
Greedy: Yes
PM: No
Figure 5: Additional qualitative examples on the POPE benchmark showing the input question, the original image, the processed perception map $\mathcal{P}$, and the perception-magnified image, arranged from left to right. The answers generated by greedy decoding and PM are compared, with correct answers highlighted in green and incorrect ones in red.
Qualitative Samples of the PM
Figure 5 presents additional qualitative examples illustrating the input question, the original image, the processed perception map $\mathcal{P}$, and the perception-magnified image. The results demonstrate that the proposed method effectively enlarges small objects relevant to the question while preserving fine details. In many cases, the VLM benefits from the perception-magnified image, as it can now recognize previously overlooked objects that were too small to be detected in the original image. This leads to more accurate predictions, as observed in the first and second rows.
Furthermore, when the VLM erroneously identifies a non-existent object as present, as shown in the third example, PM magnifies the corresponding region, enabling the model to reassess the scene with finer detail. This process helps reduce false positive hallucinations by magnifying the visual details hence improve perceptibility. However, localized magnification does not always guarantee improved visual faithfulness. As demonstrated in the fourth row, despite providing a closer view, the model now fails to recognize the cabinet in the background. This may be due to the cabinet being identified based on its structural context rather than fine-grained textures, highlighting the limitations of region-focused magnification.
Overall, these results suggest that the proposed method effectively enhances the perceptibility of visual detail on regions crucial for next-token generation while generally preserving the overall structural integrity of objects, ultimately reducing hallucinations and improving model performance.
Figure 6: A sample CLEVR-style image (left) and its corresponding magnified visual input (right). For tasks involving shape or size-related attributes, the proposed PM may introduce unintended distortions due to unregulated magnification. In this example, magnifying both the red cube and the blue sphere results in an enlarged sphere and a deformed cube, which may negatively affecting model predictions.
Limitations and Future Works
While the proposed PM method improves VLM inference by mitigating hallucinations, magnifying image regions may also introduce certain drawbacks. This section discusses key limitations of our approach.
Shape Distortion in Region Magnification.
Although region magnification allows the model to observe fine-grained visual details without altering the overall image structure, it may distort local shape properties. Specifically, in tasks requiring precise shape or length recognition, such as distinguishing letters or assessing object sizes, PM does not guarantee proportional magnification. This can introduce challenges for accurate recognition and induces hallucinations. For instance, when applied to CLEVR-style questions, as shown in Fig. 6, with the query “Please estimate how many times large is the blue sphere compared to the red cube?”, PM magnifies both objects, causing the blue sphere to appear visually larger while distorting the shape of the other two cubes. Such unintended modifications may negatively impact final predictions. Future work could explore refinements in perception map construction or post-processing techniques to achieve more structure-preserving magnification. To address the issue of distortion, several enhancements can be integrated into our framework. One potential direction involves incorporating detected bounding boxes to constrain the perception map, thereby regularizing the magnification process and preserving object shape integrity. Another approach is to group visual tokens with shared semantic representations, enabling uniform resampling across coherent regions and reducing uneven deformation within individual objects. Furthermore, for tasks where geometric fidelity is critical—such as those involving precise reasoning over shape or size—magnification can be applied adaptively at selective decoding steps rather than continuously, allowing better control over the trade-off between visual detail enhancement and structural preservation.
Efficiency.
A key limitation of our method lies in its computational overhead: applying perception-based magnification at every decoding step requires feature extraction from a modified image, which precludes the use of efficient VLM decoding techniques such as key-value (KV) caching. Nonetheless, our method remains relatively efficient when used without iterative refinement, as the perception map can be constructed with only a single additional VLM forward pass. To further mitigate runtime cost, the method can be adaptively applied only at decoding steps where visual information is critical (e.g., during object description) while being omitted for tokens with lower visual dependency, such as punctuation or reasoning-related words. A practical indicator for such adaptation could be the aggregated attention weights on visual tokens, signaling when the model is actively grounding its predictions in the visual input. Additionally, to further enhance efficiency, future work may explore speculative decoding strategies, where perception magnification is applied post hoc to refine previously decoded sequences. This deferred application may significantly reduce the frequency of image resampling, improving scalability for longer sequences without sacrificing generation quality.
Improving Compatibility.
While the proposed method can be easily applied to interleaved VLMs with visual tokens corresponding to all patch features, applying it to VLMs where the correlation between visual tokens and image regions are non-trivial, can be slightly more complicated. The method depends on certain mapping from the visual token attention weights to the attention span on the original image. Therefore, when apply PM to models where the visual tokens are constructed with not only a feature extractor, but also a token sampler or cross-attention feature extractors, PM will need some attention mapping mechanism between visual tokens and patches. However, we can resolve this problem by introducing gradient backpropagation to address the implicit relationships between the image and tokens, refining the visual attentions and construct the perception map, similar to method proposed in (Chefer, Gur, and Wolf 2021). Overall, such limitation does not reduce the contribution of this work, as the idea of adaptive-magnification can also always be used in aligned with interpretability tools designed for different VLMs. Alternatively, one could use PM in ways similar to (Yu, Yu, and Wang 2024), where any compatible model could be used as external guidance to the decoding-adaptive magnification.
Prompt for Reference-based Evaluation
[Context]
An aerial view of Diamond Head in the Hawaiian Islands.
[Question]
What is the name of this famous sight in the photo?
[Assistant 1]
The famous sight in the photo is Diamond Head.
[End of Assistant 1]
[Assistant 2]
The famous sight in the photo is the Na Pali Coast, which is a scenic coastal area on the island of Kaua’i in Hawaii.
[End of Assistant 2]
[System]
We would like to request your feedback on the performance of two AI assistants in response to the user question displayed above. The user asks the question on observing an image. For your reference, the visual content in the image is represented with a few sentences describing the image.
Please rate the helpfulness, relevance, accuracy, level of details of their responses. Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance.
Please first output a single line containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space.
In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgement.
Response:
10 2
Assistant 1 correctly identified the famous sight in the photo as Diamond Head, which is consistent with the context provided. The response is accurate, relevant, and directly answers the question with the appropriate level of detail given the context.
Assistant 2, on the other hand, incorrectly identified the sight as the Na Pali Coast, which is not mentioned in the context. This response is inaccurate and irrelevant to the question, as it does not align with the description of the image provided. Additionally, the details about the Na Pali Coast, while informative, are not applicable to the context of Diamond Head. Therefore, Assistant 2’s response is not helpful in this scenario.
Table 8: Prompt for reference-based evaluation.
GPT-Assisted Evaluation
We adopt the GPT-assisted evaluation prompts directly from prior studies. For reference-based evaluation, we follow the prompt used in (Liu et al. 2023), with an example of the prompt and corresponding evaluation result provided in Tab. 8. For vision-based evaluation, we use the prompt from (Huang et al. 2024), with an example shown in Tab. 9.
Evaluation Protocol
Each tested method is evaluated under two settings:
  • Reference-based Evaluation: GPT-4o is given the question along with a detailed textual description of the image.
  • Vision-based Evaluation: GPT-4o is given the question and the original image.
In both settings, GPT-4o scores two answers: one from the tested method and another from GPT-4 (used as the reference). Each answer is rated independently on a scale of 0 to 10. The final evaluation score is computed as:
$\text{Relative Score}=\frac{\text{Score}_{\text{method}}}{\text{Score}_{\text{GPT-4}}}$
This normalization ensures comparability across questions of varying difficulty.
Evaluation Criteria
The evaluation criteria differ slightly between the two settings:
  • In reference-based evaluation, GPT-4o assesses overall content quality, coherence, and informativeness of the response relative to the image description.
  • In vision-based evaluation, GPT-4o is explicitly instructed to focus on the accuracy of visual grounding and the level of detail present in the answer.
Prompt for Vision-based Evaluation
You are required to score the performance of two AI assistants in describing a given image. You should pay extra attention to the hallucination, which refers to the part of descriptions that are inconsistent with the image content, such as claiming the existence of something not present in the image or describing incorrectly in terms of the counts, positions, or colors of objects in the image. Please rate the responses of the assistants on a scale of 1 to 10, where a higher score indicates better performance, according to the following criteria:
1: Accuracy: whether the response is accurate with respect to the image content. Responses with fewer hallucinations should be given higher scores.
2: Detailedness: whether the response is rich in necessary details. Note that hallucinated descriptions should not count as necessary details.
Please output the scores for each criterion, containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. Following the scores, please provide an explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgement.
[Assistant 1]
The intended effect of this painting might be to create a humorous and whimsical twist on the famous Mona Lisa artwork by featuring a dog dressed in her attire, which can evoke feelings of amusement and curiosity.
[End of Assistant 1]
[Assistant 2]
The intended effect of this painting is to create a humorous and whimsical scene by combining a dog with a human-like appearance, such as a Renaissance-style painting. The dog is wearing a Renaissance-style dress and a hat, which adds a touch of humor and creativity to the image. The painting might be intended to evoke amusement and curiosity in the viewer, as it is not common to see a dog dressed in such a manner. The artist might have wanted to challenge the viewer’s expectations and create a unique, memorable piece of art.
[End of Assistant 2]
Output format:
Accuracy: <Scores of the two answers>
Reason:
Detailedness: <Scores of the two answers>
Reason:
Response:
“9”: **Accuracy: 8 9**
*Reason:**
Assistant 1 accurately describes the image as a humorous twist on the Mona Lisa featuring a dog, but lacks specific details about the attire. Assistant 2 provides a more detailed and accurate description, mentioning the Renaissance-style dress and the humorous intent, which aligns well with the image content.
*Detailedness: 6 9**
*Reason:**
Assistant 1 gives a brief description focusing on the humorous aspect but lacks detail about the attire and style. Assistant 2 offers a richer description, noting the Renaissance-style dress and hat, and elaborates on the artistic intent, providing a more comprehensive understanding of the image.”
Table 9: Prompt for vision-based evaluation, note the image is processed by GPT-4o for evaluation.